### v1

# Single-loop Algorithms for Stochastic Non-convex Optimization with Weakly-Convex Constraints

Abstract
Constrained optimization with multiple functional inequality constraints has significant applications in machine learning. This paper examines a crucial subset of such problems where both the objective and constraint functions are weakly convex. Existing methods often face limitations, including slow convergence rates or reliance on double-loop algorithmic designs. To overcome these challenges, we introduce a novel single-loop penalty-based stochastic algorithm. Following the classical exact penalty method, our approach employs a hinge-based penalty, which permits the use of a constant penalty parameter, enabling us to achieve a state-of-the-art complexity for finding an approximate Karush-Kuhn-Tucker (KKT) solution. We further extend our algorithm to address finite-sum coupled compositional objectives, which are prevalent in artificial intelligence applications, establishing improved complexity over existing approaches. Finally, we validate our method through experiments on fair learning with receiver operating characteristic (ROC) fairness constraints and continual learning with non-forgetting constraints.
1 Introduction
In this paper, we consider solving the following non-linear inequality constrained optimization problem:
$\min_{\mathbf{x}}F(\mathbf{x}),~{}~{}~{}\text{s.t. }~{}~{}h_{k}(\mathbf{x})\leq 0,\quad k=1,\ldots,m,$ (1)
where $F$ and $h_{k}$, $k=1,\dots,m$ are stochastic, weakly convex and potentially non-smooth. This general formulation captures a wide range of practical optimization problems where non-smoothness arises in the objective and constraint functions. Inequality constrained optimization arise in many applications in ML/AI, including continual learning [21], fairness-aware learning [35], multi-class Neyman-Pearson classification (mNPC) problem [27], robust learning [30], policy optimization problem for reinforcement learning [33], and distributed data center scheduling problem [39].
In optimization literature, many techniques have been studied for solving convex constrained optimization problems, including primal-dual methods [6, 29], augmented Lagrangian methods [5], penalty methods [7], and level-set methods [24]. However, the presence of non-convexity in the problem often complicates algorithmic design and convergence analysis, while non-smoothness and stochasticity add additional layers of difficulty.
Several recent studies have explored stochastic non-convex constrained optimization under various assumptions. Ma et al. [27] and Boob et al. [6] initiated the study of stochastic optimization with weakly convex and non-smooth objective and constraint functions. Inspired by proximal-point methods for weakly-convex unconstrained optimization [10], they have proposed a double-loop algorithm, where a sequence of convex quadratically regularized subproblems is solved approximately. They have derived the state-of-the-art complexity in the order of $O(1/\epsilon^{6})$ for finding a nearly $\epsilon$-KKT solution under a uniform Slater’s condition and a strong feasibility condition, respectively.
Recently, Alacaoglu and Wright [1], Li et al. [21], Huang and Lin [17] have proposed single-loop algorithms for stochastic non-convex constrained optimization. However, these results either are restricted to smooth problems with equality constraints [1] or suffer from an iteration complexity higher than $O(1/\epsilon^{6})$ [21, 17]. This raises an intriguing question: “Can we design a single-loop method for solving non-smooth weakly-convex stochastic constrained optimization problems to match the state-of-the-art complexity of $O(1/\epsilon^{6})$ for finding a nearly $\epsilon$-KKT solution?"
Table 1: Comparison with existing works. We consider that both objective function and constraint function oracles are stochastic. In the Constraint column, $h_{k}$ means multiple constraints, $k$ from 1 to $m$. In Convexity column, NC means Non-convex, WC means Weakly Convex, C means Convex. The monotonicity column means non-decreasing. The last column Complexity means the complexity for finding an (nearly) $\epsilon$-KKT point of (1).
Reference Loop Objective Constraints Smoothness Convexity Monotonicity Complexity
[1] Single Loop $F(\cdot)$ $h(\cdot)=0$ $F,h$ NC ($F,h$) none $\tilde{\mathcal{O}}(\epsilon^{-5})$
[23] Double Loop $F(\cdot)+g(\cdot)$ $h(\cdot)=0$ $F,h$ C ($g$), NC ($F,h$) none $\tilde{\mathcal{O}}(\epsilon^{-5})$
[27] Double Loop $F(\cdot)$ $h_{k}(\cdot)\leq 0$ none WC ($F,h_{k}$) none $\mathcal{O}(\epsilon^{-6})$
[6] Double Loop $F(\cdot)$ $h_{k}(\cdot)\leq 0$ none WC ($F,h_{k}$) none $\mathcal{O}(\epsilon^{-6})$
[17] Single Loop $F(\cdot)$ $h(\cdot)\leq 0$ none WC ($F$), C ($h$) none $\tilde{\mathcal{O}}(\epsilon^{-8})$
[21] Single Loop $\sum_{i}f_{i}(g_{i}(\cdot))$ $h_{k}(\cdot)\leq 0$ $f_{i},g_{i},h_{k}$ NC ($f_{i},g_{i},h_{k}$) none $\mathcal{O}(\epsilon^{-7})$
Ours Single Loop $F(\cdot)$ $h_{k}(\cdot)\leq 0$ none WC($F,h_{k}$) none $\mathcal{O}(\epsilon^{-6})$
Ours Single Loop $\sum_{i}f_{i}(g_{i}(\cdot))$ $h_{k}(\cdot)\leq 0$ none WC ($f_{i}$, $g_{i}$, $h_{k}$) $f_{i}$ $\mathcal{O}({\epsilon^{-6}})$
Ours Single Loop $\sum_{i}f_{i}(g_{i}(\cdot))$ $h_{k}(\cdot)\leq 0$ $f_{i},g_{i}$ WC ($h_{k}$) none $\mathcal{O}(\epsilon^{-6})$
We answer the question in the affirmative. We propose a novel algorithm based on the well-studied exact penalty method [42, 13]. Different from Alacaoglu and Wright [1], Li et al. [21] who used a quadratic or squared hinge penalty, we construct a penalty of the constraints using the non-smooth hinge function, i.e., by adding $\frac{1}{m}\sum_{k=1}^{m}\beta[h_{k}(\mathbf{x})]_{+}$ into the objective, where $\beta$ is a penalty parameter. Despite being non-smooth, the benefits of this penalty function include (i) it preserves the weak convexity of in the penalty; (ii) it allows using a constant $\beta$ as opposed to a large $\beta$ required by the squared hinge penalty [21]; (iii) it introduces the structure of finite-sum coupled compositional optimization (FCCO), enabling the use of existing FCCO techniques in the algorithmic design. Consequently, this approach establishes state-of-the-art complexity for achieving a nearly $\epsilon$-KKT solution, under a regularity assumption similar to those in [21, 1]. We summarize our contributions as follows.
  • We propose a non-smooth, hinge-based exact penalty method for tackling non-convex constrained optimization with weakly convex objective and constraint functions. We derive a theorem to guarantee that a nearly $\epsilon$-stationary solution of the penalized objective is a nearly $\epsilon$-KKT solution of the original problem under a regularity condition of the constraints.
  • We develop algorithms based on FCCO to tackle a non-smooth weakly convex objective whose unbiased stochastic gradient is available and a structured non-smooth weakly-convex objective that is of the FCCO form. We derive a state-of-the-art complexity of $\mathcal{O}(\epsilon^{-6})$ for finding a nearly $\epsilon$-KKT solution.
  • We conduct experiments on two applications in machine learning: learning with ROC fairness constraints and continual learning with non-forgetting constraints. The effectiveness of our algorithms are demonstrated in comparison with an existing method based on squared hinge penalty function and a double-loop method.
2 Related Work
Despite their long history of study, recent literature on optimization with non-convex constraints has introduced new techniques and theories for the penalty method [25], Lagrangian methods [31, 34, 23, 22], the sequential quadratic programming methods [3, 4, 8, 28] and trust region method [14]. These works mainly focused on the problems with smooth objective and constraint functions. On the contrary, based on Moreau envelopes, Ma et al. [27], Boob et al. [6], Huang and Lin [17], Jia and Grimmer [19] have developed algorithms and theories for problems with non-smooth objective and constraint functions. However, these methods cannot be applied when the objective function has a compositional structure. The method by Li et al. [21] can be applied when the objective function has a coupled compositional structure, but it requires the objective and constraint functions to be smooth. Compared to their work, we relax the smoothness assumption, but assume the monotonicity of the outer function in the compositional structure.
Alacaoglu and Wright [1] studies a single-loop quadratic penalty method for stochastic smooth nonconvex equality constrained optimization under a generalized full-rank assumption on the Jacobian of the constraints. Under a similar assumption, Li et al. [23] proposed a double-loop inexact augmented Lagrangian method using the momentum-based variance-reduced proximal stochastic gradient algorithm as a subroutine. Both Alacaoglu and Wright [1] and Li et al. [23] consider only smooth equality constraints and both achieved a complexity of $\mathcal{O}(1/\epsilon^{5})$.  [^1] Complexity lower than $\mathcal{O}(1/\epsilon^{5})$ can be achieved by Alacaoglu and Wright [1] when the constraints are deterministic. Ma et al. [27] and Boob et al. [6] studied weakly convex non-smooth optimization with inequality constraints. They proposed double-loop methods, where a sequence of proximal-point subproblems is solved approximately by a stochastic subgradient method [39] or a constraint extrapolation method [6]. Under a uniform Slater’s condition and a strong feasibility assumption, respectively, Ma et al. [27] and Boob et al. [6] establish a complexity of $\mathcal{O}(1/\epsilon^{6})$. Under the uniform Slater’s condition, Huang and Lin [17] proposed a single-loop switching subgradient method for non-smooth weakly convex constrained optimization problems, but the complexity of their method is $\mathcal{O}(1/\epsilon^{8})$ when the constraints are stochastic.
Motivated by applications from continual learning with non-forgetting constraints, Li et al. [21] introduced a single-loop squared hinge penalty method for a FCCO [36] problem, where the objective function takes the form of $\sum_{i}f_{i}(g_{i}(\cdot))$ and the constraints are non-convex and smooth. They establish a complexity of $\mathcal{O}(1/\epsilon^{7})$ under the generalized full-rank assumption on the Jacobian of constraints. In our work, we consider a FCCO problem with nonconvex but unnecessarily smooth inequality constraints. Following the traditional exact penalty method, we first add a hinge penalty of the constraints to the objective function, and solve the unconstrained FCCO using a single-loop algorithm. We show that our method improves the complexity to $\mathcal{O}(1/\epsilon^{6})$ in two cases: (1) $f_{i}$, $g_{i}$ and $h_{k}$ are non-smooth weakly convex and $f_{i}$ is monotone non-decreasing, and (2) $f_{i}$ and $g_{i}$ are smooth nonconvex but $h_{k}$ is non-smooth weakly convex. Note that the second case includes the special case where $f_{i}$, $g_{i}$ and $h_{k}$ are smooth non-convex. A comparison with prior results in presented in Table 1.
Finally, we would like to acknowledge a concurrent work by [26]  [^2] The first version of the present work [37] was submitted on 01/30/2025.. Their study also explores a penalty method for problems with weakly convex objective and constraint function. The key differences between our work and theirs are: (i) they consider only one constraint function, while we tackle multiple constraint functions with stochastic sampling; (ii) they do not directly use the hinge penalty function, but instead adopt a smoothed version of it; and (iii) they employ the SPIDER/SARAH technique to track the constraint function value, which requires periodically using large batch sizes. In contrast, we track multiple constraint function values by employing the MSVR technique [20] using only a constant batch size.
3 Notations and Preliminaries
[…]
We say $\psi$ is $\mu$-strongly convex ($\mu\geq 0$) if $\psi(\mathbf{x})\geq\psi(\mathbf{x}^{\prime})+\langle\zeta,\mathbf{x}-\mathbf{x}^{\prime}\rangle+\frac{\mu}{2}\|\mathbf{x}-\mathbf{x}^{\prime}\|^{2}$ for any $(\mathbf{x},\mathbf{x}^{\prime})$ and any $\zeta\in\partial\psi(\mathbf{x}^{\prime})$. We say $\psi$ is $\rho$-weakly convex ($\rho>0$) if $\psi(\mathbf{x})\geq\psi(\mathbf{x}^{\prime})+\langle\zeta,\mathbf{x}-\mathbf{x}^{\prime}\rangle-\frac{\rho}{2}\|\mathbf{x}-\mathbf{x}^{\prime}\|^{2}$ for any $(\mathbf{x},\mathbf{x}^{\prime})$ and any $\zeta\in\partial\psi(\mathbf{x}^{\prime}).$ We say $\psi$ is $L$-smooth ($L\geq 0$) if $|\psi(\mathbf{x})-\psi(\mathbf{x}^{\prime})+\langle\nabla\psi(\mathbf{x}^{\prime}),\mathbf{x}-\mathbf{x}^{\prime}\rangle|\leq\frac{L}{2}\|\mathbf{x}-\mathbf{x}^{\prime}\|^{2}$ for any $(\mathbf{x},\mathbf{x}^{\prime})$. For simplicity, we abuse $\partial\psi(x)$ to denote one subgradient from the corresponding subgradient set when no confusion could be caused.
Throughout the paper, we make the following assumptions on (1)
Assumption 3.1.
(a) $F$ is $\rho_{0}$-weakly convex and $L_{F}$-Lipschitz continuous; (b) $h_{k}(\cdot)$ is $\rho_{1}$-weakly convex and $L_{h}$-Lipschitz continuous for $k=1,\dots,m.$
For a non-convex optimization problem, finding a globally optimal solution is intractable. Instead, a Karush-Kuhn-Tucker (KKT) solution is of interest, which is an extension of a stationary solution of an unconstrained non-convex optimization problem. A solution $\mathbf{x}$ is a KKT solution to (1) if there exist $\boldsymbol{\lambda}=(\lambda_{1},\dots,\lambda_{m})^{\top}\in\mathbb{R}_{+}^{m}$ such that $0\in\partial F(\mathbf{x})+\sum_{k=1}^{m}\lambda_{k}\partial h(\mathbf{x})$, $h_{k}(\mathbf{x})\leq 0,\forall k$ and $\lambda_{k}h_{k}(\mathbf{x})=0$ $\forall k$. For non-asymptotic analysis, one might consider finding an $\epsilon$-KKT solution as defined below.
Definition 3.2.
A solution $\mathbf{x}$ is an $\epsilon$-KKT solution to (1) if there exist $\boldsymbol{\lambda}=(\lambda_{1},\dots,\lambda_{m})^{\top}\in\mathbb{R}_{+}^{m}$ such that $\text{dist}(0,\partial F(\mathbf{x})+\sum_{k=1}^{m}\lambda_{k}\partial h(\mathbf{x}))\leq\epsilon$, $h_{k}(\mathbf{x})\leq\epsilon,\forall k$, and $\lambda_{k}h_{k}(\mathbf{x})\leq\epsilon$ if $h_{k}(\mathbf{x})\geq 0$ or $\lambda_{k}=0$ if $h_{k}(\mathbf{x})<0$ for $k=1,\dots,m$.
However, since the objective and the constraint functions are non-smooth, finding an $\epsilon$-KKT solution is not tractable, even the constraint functions are absent. For example, if $F(x)=|x|$ finding $\epsilon$-stationary solution is infeasible unless we find the optimal solution $x=0$. To address this challenge, we follow many previous works on weakly convex optimization [10] and consider finding a nearly $\epsilon$-KKT solution defined below.
Definition 3.3.
A solution $\mathbf{x}$ is a nearly $\epsilon$-KKT solution to (1) if there exist $\bar{\mathbf{x}}$ and $\boldsymbol{\lambda}=(\lambda_{1},\dots,\lambda_{m})^{\top}\in\mathbb{R}_{+}^{m}$ such that (i) $\|\mathbf{x}-\bar{\mathbf{x}}\|\leq O(\epsilon)$, $\text{dist}(0,\partial F(\bar{\mathbf{x}})+\sum_{k=1}^{m}\lambda_{k}\partial h(\bar{\mathbf{x}}))\leq\epsilon$, (ii) $h_{k}(\bar{\mathbf{x}})\leq\epsilon,\forall k$, and (iii) $\lambda_{k}h_{k}(\bar{\mathbf{x}})\leq\epsilon$ if $h_{k}(\bar{\mathbf{x}})\geq 0$ or $\lambda_{k}=0$ if $h_{k}(\bar{\mathbf{x}})<0$ for $k=1,\dots,m$.
In order to develop our analysis, we introduce the Moreau envelope of a $\rho$-weakly convex function $\phi$:
$\phi_{\theta}(\mathbf{x}):=\min_{\mathbf{y}}\left\{\phi(\mathbf{y})+\frac{1}{2\theta}\|\mathbf{y}-\mathbf{x}\|^{2}\right\}$
where $\theta<\rho^{-1}$.
The Moreau envelope is an implicit smoothing of the original problem. The following lemma follows from standard result [9].
Lemma 3.4.
Given a $\rho$ weakly convex function $\phi$ and $\theta<(\rho)^{-1}$, then the envelope $\phi_{\theta}$ is smooth with gradient given by $\nabla\phi_{\theta}(\mathbf{x})=\theta^{-1}(\mathbf{x}-\text{prox}_{\theta\phi}(\mathbf{x}))$, where
$\text{prox}_{\theta\phi}(\mathbf{x}):=\arg\min_{\mathbf{y}}\left\{\phi(\mathbf{y})+\frac{1}{2\theta}\|\mathbf{y}-\mathbf{x}\|^{2}\right\}.$ (2)
The smoothness constant of $\phi_{\theta}$ is $\frac{2-\theta\rho}{\theta(1-\theta\rho)}$. In addition, $\text{dist}(0,\partial\phi(\bar{\mathbf{x}}))\leq\|\nabla\phi_{\theta}(\mathbf{x})\|$.
Hence, if we find $\mathbf{x}$ such that $\|\nabla\phi_{\theta}(\mathbf{x})\|\leq\epsilon$, then we can say that $\mathbf{x}$ is close to a point $\bar{\mathbf{x}}$ that is $\epsilon$-stationary. We call $\mathbf{x}$ a nearly $\epsilon$-stationary solution of $\min_{\mathbf{x}}\phi(\mathbf{x})$.
4 A Hinge Exact Penalty Method and Theory
The key idea of the hinge-based exact penalty method is to solve the following unconstrained minimization [42, 13]:
$\min_{\mathbf{x}}\Phi(\mathbf{x}):=F(\mathbf{x})+\frac{\beta}{m}\sum_{k=1}^{m}[h_{k}(\mathbf{x})]_{+},$ (3)
where $[\cdot]_{+}:=\max\{0,\cdot\}$ is the hinge function, and $\beta>0$ is a penalty parameter. For clarity of notation, we define
$h_{k}^{+}(\mathbf{x})=[h_{k}(\mathbf{x})]_{+}.$
Let $\partial[h_{k}(x)]_{+}\in[0,1]$ be the subdifferential of $[z]_{+}$ at $z=h_{k}(x)$ (different from $\partial h_{k}^{+}(\mathbf{x})$).
Lemma 4.1.
Under Assumption 3.1, $\Phi(\mathbf{x})$ in (3) is $C$-weakly convex and $L$-Lipschitz, where $C=\rho_{0}+\beta\rho_{1}$ and $L=L_{F}+\beta L_{h}$.
Our hinge exact penalty method is to employ a stochastic algorithm for finding an $\epsilon$-stationary solution to $\Phi_{\theta}(\mathbf{x})$ for some $\theta<1/C$. We establish a theorem below to guarantee that such $\mathbf{x}$ is a nearly $\epsilon$-KKT solution to (1).
Theorem 4.2.
Assume that there exists $\delta>0$ such that
$\text{dist}\left(0,\frac{1}{m}\sum_{k=1}^{m}\partial h_{k}^{+}(\mathbf{x}^{\prime})\right)\geq\delta,\forall\mathbf{x}^{\prime}\in\mathcal{V}$ (4)
where $\mathcal{V}=\{\mathbf{x}:\max_{k}h_{k}(\mathbf{x})>0\}.$ If $\mathbf{x}$ is an $\epsilon$-stationary solution to $\Phi_{\theta}(\mathbf{x})$ for some $\theta<1/C$ and a constant $\beta>(\epsilon+L_{F})/\delta$, then $\mathbf{x}$ is a nearly $\epsilon$-KKT solution to the original problem (1).
Proof.
Let $\bar{\mathbf{x}}=\text{prox}_{\theta\Phi}(\mathbf{x})$. By optimality of $\bar{\mathbf{x}}$, we have
$0\in\partial F(\bar{\mathbf{x}})+\frac{\beta}{m}\sum_{k=1}^{m}\partial h_{k}^{+}(\bar{\mathbf{x}})+(\bar{\mathbf{x}}-{\mathbf{x}})/\theta.$
Since $\|\nabla\Phi_{\theta}(\mathbf{x})\|\leq\epsilon$, then we have $\|\bar{\mathbf{x}}-\mathbf{x}\|=\theta\|\nabla\Phi_{\theta}(\mathbf{x})\|\leq\theta\epsilon$. Then we have
$\text{dist}\left(0,\partial F(\bar{\mathbf{x}})+\frac{\beta}{m}\sum_{k=1}^{m}\partial h_{k}^{+}(\bar{\mathbf{x}})\right)\leq\|(\bar{\mathbf{x}}-{\mathbf{x}})/\theta\|\leq\epsilon.$
Since $\partial h_{k}^{+}(\bar{\mathbf{x}})=\xi_{k}\partial h_{k}(\bar{\mathbf{x}})$ (see [2, Corollary 16.72]), where
$\xi_{k}=\begin{cases}1&\text{if }h_{k}(\bar{\mathbf{x}})>0,\\ [0,1]&\text{if }h_{k}(\bar{\mathbf{x}})=0,\\ 0&\text{if }h_{k}(\bar{\mathbf{x}})<0,\end{cases}\quad\in\partial[h_{k}(\bar{\mathbf{x}})]_{+},$
there exists $\lambda_{k}\in\frac{\beta\xi_{k}}{m}\geq 0,\forall k$ such that
$\text{dist}\left(0,\partial F(\bar{\mathbf{x}})+\sum_{k=1}^{m}\lambda_{k}\partial h_{k}(\bar{\mathbf{x}})\right)\leq\epsilon.$
Thus, we prove condition (i) in Definition 3.3. Next, let us prove condition (ii). We argue that $\max_{k}h_{k}(\bar{\mathbf{x}})\leq 0$. Suppose this does not hold, i.e., $\max_{k}h_{k}(\bar{\mathbf{x}})>0$, we will derive a contradiction. Since $\exists\mathbf{v}\in\partial F(\bar{\mathbf{x}})$ such that
$\epsilon\geq\text{dist}\left(0,\mathbf{v}+\frac{\beta}{m}\sum_{k=1}^{m}\partial h_{k}^{+}(\bar{\mathbf{x}})\right)\geq\text{dist}\left(0,\frac{\beta}{m}\sum_{k=1}^{m}\partial h_{k}^{+}(\bar{\mathbf{x}})\right)-\|\mathbf{v}\|\geq\beta\delta-L_{F},$ (5)
which is a contradiction to the assumption that $\beta>(\epsilon+L_{F})/\delta$. Thus, $\max_{k}h_{k}(\bar{\mathbf{x}})\leq 0$. This proves condition (ii). The last condition (iii) holds because: $\lambda_{k}=\frac{\beta\xi_{k}}{m}$, which is zero if $h_{k}(\bar{\mathbf{x}})<0$; and $\lambda_{k}h_{k}(\bar{\mathbf{x}})\leq 0$. ∎
4.1 Sufficient Conditions for Eq. (4)
A key assumption in Theorem 4.2 is the condition in (4). Next, we provide some sufficient conditions. For simplicity of exposition and comparison with prior works, we consider differentiable constraint functions.
First, we consider a simple setting where there is only one constraint $h(\mathbf{x})\leq 0$. The lemma below states two conditions sufficient for proving the condition (4). The proof of Lemma 4.3 and 4.4 are presented in Appendix B.
Lemma 4.3.
If there exists $\mu>0,c>0$ such that (i) $h(\mathbf{x})-\min_{\mathbf{y}}h(\mathbf{y})\leq\frac{1}{2\mu}\|\nabla h(\mathbf{x})\|^{2}$ for any $\mathbf{x}$ satisfying $h(\mathbf{x})>0$; and (ii) $\min_{\mathbf{y}}h(\mathbf{y})\leq-c$, then condition (4) holds for $\delta=\sqrt{2c\mu}$.
Remark: Note that the condition (i) in the above lemma is the PL condition of $h$ but only for points violating the constraint. We can easily come up a function $h$ to satisfy the two conditions in the lemma. Let us consider $h(x)=|x^{2}-1|-1$. Then $\nabla h(0)=0$, which means $x=0$ is a stationary point but not global optimum of $h$. The global optimal are $x\in\{1,-1\}$ and $\min_{y}h(y)=-1$, and hence condition (ii) holds. Any $x$ such that $h(x)>0$ would satisfy $x>\sqrt{2}$ or $x<-\sqrt{2}$. For these $x$, we have $\nabla h(x)=2x$ for $x>\sqrt{2}$ or $\nabla h(x)=-2x$ for $x<-\sqrt{2}$. Then for $|x|>\sqrt{2}$, $h(x)-\min_{y}h(y)=x^{2}-1\leq\frac{|\nabla h(x)|^{2}}{2\mu}$ holds for $\mu=2$. Hence condition (i) holds.
Next, we present a sufficient condition for (4) for multiple constraints. The following lemma shows that the full rank assumption of the Jacobian matrix of $H(\mathbf{x})=(h_{1}(\mathbf{x}),\ldots h_{m}(\mathbf{x}))^{\top}$ at $x$ that violate constraints is a sufficient condition.
Lemma 4.4.
If $\nabla H(\mathbf{x})$ is full rank for $\max_{k}h_{k}(\mathbf{x})>0$, i.e., $\lambda_{min}(\nabla H(\mathbf{x}))\geq\sigma>0,\forall\mathbf{x}\text{ such that }\max_{k}h_{k}(\mathbf{x})>0$, where $\lambda_{\min}(\cdot)$ denotes the minimum eigen-value of a matrix, then the condition (4) holds for $\delta=\sigma/m$.
We refer to the above condition as Full-Rank-at-Violating-Points condition (FRVP). We note that the FRVP condition has been used as a sufficient condition in [21] for ensuring
$\|\nabla H(\mathbf{x})^{\top}[H(\mathbf{x})]_{+}\|\geq\delta\|[H(\mathbf{x})]_{+}\|,\forall\mathbf{x},$ (6)
which is central to proving the convergence of their squared-hinge penalty method for solving inequality constrained smooth optimization problems: $\min_{\mathbf{x}}F(\mathbf{x}),\text{ s.t. }~{}h_{k}(\mathbf{x})\leq 0,k=1,\ldots,m$.
A stronger condition for equality constrained optimization with $h_{k}(\mathbf{x})=0,k=1,\ldots,m$ was imposed in [1], where they assumed that
$\|\nabla H(\mathbf{x})^{\top}H(\mathbf{x})\|\geq\delta\|H(\mathbf{x})\|,\forall\mathbf{x}.$ (7)
To ensure this, their method requires the full-rank Jacobian assumption to hold at all points, whereas our approach only requires it at constraint-violating points. In our experiments, we verified that the condition in Lemma 4.4 holds on the trajectory of the algorithm. Detailed results are provided in Table 2 in Appendix A.
4.2 Issue with squared-hinge penalty
Let us discuss why using the squared-hinge penalty function as in [21]. Using a squared-hinge penalty function gives the following objective:
$\Phi(\mathbf{x})=F(\mathbf{x})+\frac{\beta}{m}\sum_{k=1}^{m}[h_{k}(\mathbf{x})]^{2}_{+}.$ (8)
Note that this objective is also weakly convex with a weak-convexity parameter given by $\rho_{0}+\beta\rho_{1}$. However, the key issue of using this approach is that $\beta$ needs to be very large to find an (nearly) $\epsilon$-KKT solution. Using the same argument as in the proof of Theorem 4.2 for showing $\max_{k}h_{k}(\bar{\mathbf{x}})\leq 0$, the inequality (5) becomes:
$\epsilon\geq\left\|\mathbf{v}+\frac{\beta}{m}\sum_{k=1}^{m}[h_{k}(\bar{\mathbf{x}})]_{+}\nabla h_{k}(\bar{\mathbf{x}})\right\|\geq\left\|\frac{\beta}{m}\sum_{k=1}^{m}[h_{k}(\bar{\mathbf{x}})]_{+}\nabla h_{k}(\bar{\mathbf{x}})\right\|-\|\mathbf{v}\|\geq\frac{\beta\delta}{m}\|[H(\bar{\mathbf{x}})]_{+}\|-L_{F},$
where the last step follows the assumption in (6). Since $[H(\bar{\mathbf{x}})]_{+}$ could be very small, in order to derive a contradiction we need to set $\beta$ to be very large. In [21], it was set to be the order of $O(1/\epsilon)$. As a result, $\Phi(\cdot)$ will have large weak-convexity and Lipschitz parameters of the order $O(1/\epsilon)$. This will lead to a complexity of $O(1/\epsilon^{12})$ if we follow existing studies to solve the problem in (8[16] without the smoothness of $F$ and $h_{k}$, which will be discussed in next section.
4.3 Issue of solving an equivalent equality constrained optimization
Does converting the inequality constrained problem into an equality constrained problem work? In general this answer is no. One may consider the following equivalent equality constrained problem:
$\min_{\mathbf{x},s_{1},\ldots,s_{m}}F(\mathbf{x}),~{}~{}~{}\text{s.t. }~{}~{}h_{k}(\mathbf{x})+s_{k}^{2}=0,k=1,\ldots,m$ (9)
where $s_{1},\ldots,s_{m}$ are introduced dummy variables. A solution $\mathbf{x}^{*}$ is a KKT solution to above if there exists $\boldsymbol{\lambda}=(\lambda_{1},\ldots,\lambda_{m})^{\top}\in\mathbb{R}^{m}$ such that $0\in\partial F(\mathbf{x}^{*})+\sum_{k=1}^{m}\lambda_{k}\partial h_{k}(\mathbf{x}^{*}),$ $\lambda_{k}s_{k}=0,\forall k$, and $h_{k}(\mathbf{x}^{*})+s_{k}^{2}=0$ for $\forall{k}=1,\ldots,m$. However, this does not provide a guarantee that $\mathbf{x}_{*}$ is a KKT condition to the original constrained optimization problem as $\boldsymbol{\lambda}$ is not necessarily non-negative. For instance, consider $\min x\text{ s.t. }-1\leq x\leq 1$. After adding slack variables, we have $\min x\text{ s.t. }x+s_{1}^{2}=1,-x+s_{2}^{2}=1$. However, $(x,s_{1},s_{2})=(1,0,\sqrt{2})$ becomes a stationary point with multipliers $\lambda=(-1,0)$, but it is not a stationary point in the original problem.
While several studies [15, 11] have employed stronger second-order conditions to ensure that a KKT solution of the equality-constrained problem (9) can be converted into one for the original inequality-constrained problem, these conditions generally do not hold for the problems we consider. Moreover, when these second-order conditions fail to be satisfied, the approach of solving the equality-constrained problem (9) becomes ineffective. Our solutions will not suffer from this issue.
5 Stochastic Algorithms
In this section, we consider stochastic algorithms for solving (3) under different settings for the objective function $F(\mathbf{x})$. The first setting is the following:
[…]
which has been considered in [27, 17, 6]. The second setting is:
[…]
which has been considered in [21].
Without loss of generality, we assume the constrained functions are only accessed through stochastic samples.
Assumption 5.1.
For any $\mathbf{x}$, assume there exists $h_{k}(\mathbf{x},\xi)$ and $\partial h_{k}(\mathbf{x},\xi)$, where $\xi$ is a random variable, such that (a) $\mathbb{E}[h_{k}(\mathbf{x},\xi)]=h_{k}(\mathbf{x})$, and $\mathbb{E}[\partial h_{k}(\mathbf{x},\xi)]\in\partial h_{k}(\mathbf{x})$; (b) $\mathbb{E}_{\xi}[|h_{k}(\mathbf{x},\xi)-h_{k}(\mathbf{y},\xi)|^{2}]\leq L_{h}^{2}\|\mathbf{x}-\mathbf{y}\|^{2}$; (c) $\mathbb{E}[|h_{k}(\mathbf{x},\xi)-h_{k}(\mathbf{x})|^{2}]\leq\sigma_{h}^{2}$.
Below, we let $h(\mathbf{x},\mathcal{B})=\frac{1}{|\mathcal{B}|}\sum_{\xi\in\mathcal{B}}h(\mathbf{x},\xi)$ with a mini-batch of samples $\mathcal{B}$.
5.1 Setting I
We make the following assumption regarding $f$.
Assumption 5.2.
For any $\mathbf{x}$, $\mathbb{E}_{\zeta}[\|\partial F(\mathbf{x})-\partial f(\mathbf{x},\zeta)\|^{2}]\leq\sigma_{f}^{2}$.
The key to our algorithm design is to notice that the hinge-based exact penalty function $\frac{\beta}{m}\sum_{k=1}^{m}[h_{k}(\mathbf{x})]_{+}$ is a special case of non-smooth weakly-convex finite-sum coupled compositional objective as considered in [16], where the inner functions are $h_{k}(\mathbf{x})$ and the outer function is the hinge function that is monotone and convex. Hence, we can directly utilize their technique for computing a stochastic gradient estimator of the penalty function.
The subgradient of the penalty function is given by $\frac{1}{m}\sum_{k=1}^{m}\beta[h_{k}(\mathbf{x})]^{\prime}_{+}\partial h_{k}(\mathbf{x})$, where $[h_{k}(\mathbf{x})]^{\prime}_{+}\in\partial[h_{k}(\mathbf{x})]_{+}$ is a subgradient of the hinge function $[z]_{+}$ evaluated at $z=h_{k}(\mathbf{x})$. However, computing $h_{k}(\mathbf{x})$ and $\partial h_{k}(\mathbf{x})$ is prohibited due to its stochastic nature or depending on many samples. To this end, we need an estimator for $h_{k}(\mathbf{x})$ and $\partial h_{k}(\mathbf{x})$. $\partial h_{k}(\mathbf{x})$ can be simply estimated by its mini-batch estimator. The challenge is to estimate $h_{k}(\mathbf{x})$ using mini-batch samples. The naive approach that estimates $h_{k}(\mathbf{x})$ by its mini-batch estimator does not yield a convergence guarantee for compositional optimization. To this end, we use the variance reduction technique MSVR for estimating constraint functions $h_{k}(\mathbf{x}),\forall k$ [20], which maintains and updates an estimator $u_{2,k}$ for each constraint function $h_{k}$.
In particular, at the $t$-th iteration, we construct a random mini-batch $\mathcal{B}^{t}_{c}\subset\{1,\ldots,m\}$, then we update the estimators of $h_{k}(\mathbf{x}_{t}),k\in\mathcal{B}^{t}_{c}$ by the following:
$u_{2,k}^{t+1}=(1-\gamma_{2})u_{2,k}^{t}+\gamma_{2}{h_{k}}(\mathbf{x}_{t};\mathcal{B}_{2,k}^{t})+\gamma_{2}^{\prime}({h_{k}}(\mathbf{x}_{t};\mathcal{B}_{2,k}^{t})-{h_{k}}(\mathbf{x}_{t-1};\mathcal{B}_{2,k}^{t})),$ (12)
where $\gamma_{2}\in(0,1),\gamma_{2}^{\prime}=\frac{m-|\mathcal{B}_{c}|}{|\mathcal{B}_{c}|(1-\gamma_{2})}+1-\gamma_{2}$, and $\mathcal{B}^{t}_{2,k}$ is a mini-batch of samples for estimating $h_{k}(\mathbf{x}_{t})$. Then the gradient of the penalty function at $\mathbf{x}_{t}$ can be estimated by
$G_{2}^{t}=\frac{\beta}{|\mathcal{B}^{t}_{c}|}\sum_{k\in\mathcal{B}^{t}_{c}}\partial h_{k}(\mathbf{x}_{t};\mathcal{B}_{2,k}^{t})[u_{2,k}^{t}]^{\prime}_{+},$ (13)
where $[u_{2,k}^{t}]^{\prime}_{+}$ is the derivative of the hinge function at $u_{2,k}^{t}$. The subgradient of $F(\mathbf{x}_{t})$ can be simply computed by
$G_{1}^{t}=\frac{1}{|\mathcal{B}^{t}|}\sum_{\zeta\in\mathcal{B}^{t}}\partial f(\mathbf{x}_{t};\zeta),$ (14)
where $\mathcal{B}^{t}$ is a mini-batch of samples. Then, we can update $\mathbf{x}_{t+1}$ by using SGD with $G_{1}^{t}+G_{2}^{t}$ as the stochastic gradient estimator. We present the key steps in Algorithm 1. The convergence of the algorithm for finding an $\epsilon$-stationary solution $\Phi_{\lambda}(x)$ is given below. The proof of the theorem is presented in Appendix C.
Algorithm 1 Algorithm for solving (3) under Setting I
1:  Initialization: choose $\mathbf{x}_{0},\beta,\gamma_{2}$ and $\eta$.
2:  for $t=0$ to $T-1$ do
3:     Sample $\mathcal{B}^{t}_{c}\in\{1,\ldots,m\}$
4:     Compute $G_{1}^{t}$ in Eqn (14)
5:     for  each $k\in\mathcal{B}^{t}_{c}$ do
6:        Sample a mini-batch $\mathcal{B}_{2,k}^{t}$
7:        Update the estimator $u_{2,k}^{t+1}$ by Eqn (12)
8:     end for
9:     Let $u_{2,k}^{t+1}=u_{2,k}^{t},k\notin\mathcal{B}^{t}_{c}$
10:     Compute $G_{2}^{t}$ as in Eqn (13)
11:     Update $\mathbf{x}_{t+1}=\mathbf{x}_{t}-\eta(G_{1}^{t}+G_{2}^{t})$
12:  end for
Theorem 5.3.
Consider the problem (3) under setting I and Assumptions 3.1 5.1 and 5.2. Let $\gamma_{2}=\mathcal{O}(\frac{|\mathcal{B}_{2,k}|\epsilon^{4}}{\beta^{4}})$, $\eta=\mathcal{O}(\frac{|\mathcal{B}_{c}||\mathcal{B}_{2,k}|^{1/2}\epsilon^{4}}{\beta^{5}m})$, after $T=\mathcal{O}(\frac{\beta^{6}m}{|\mathcal{B}_{c}||\mathcal{B}_{2,k}|^{1/2}\epsilon^{6}})$ iterations, the Algorithm 1 satisfies $\mathbb{E}[\|\nabla\Phi_{\theta}(\mathbf{x}_{\hat{t}})\|]\leq\epsilon$ from some $\theta=O(1/(\rho_{0}+\rho_{1}\beta))$, where $\hat{t}$ selected uniformly at random from $\{1,\ldots,T\}$.
Remark: Combining the above result and that in Theorem 4.2, Algorithm 1 can find a nearly $\epsilon$-KKT solution to the original problem (1) with $T=O(\frac{1}{\epsilon^{6}})$ iterations, which matches the state-of-the-art complexity of double-loop algorithms [27, 6]. It is better than the complexity of $O(1/\epsilon^{8})$ of the single-loop algorithm proposed in [17].
5.2 Setting II
Now, let us consider the second setting where $F(\mathbf{x})=\frac{1}{n}\sum_{i=1}^{n}f_{i}(\mathbb{E}_{\zeta}[g_{i}(\mathbf{x},\zeta)])$. Indeed, this objective is of the same form of the penalty function, which is a coupled compositional function. Let $g_{i}(\mathbf{x})=\mathbb{E}_{\zeta}[g_{i}(\mathbf{x},\zeta)]$. We make the following assumption regarding $f_{i},g_{i}$.
Assumption 5.4.
Assume $f_{i}$ is deterministic and $L_{f}$-Lipchitz continuous. For any $\mathbf{x},\mathbf{y}$, $\mathbb{E}_{\zeta}[|g_{i}(\mathbf{x},\zeta)-g_{i}(\mathbf{y},\zeta)|^{2}]\leq L_{g}^{2}\|\mathbf{x}-\mathbf{y}\|^{2}$, $\mathbb{E}[|g_{i}(\mathbf{x},\zeta)-g_{i}(\mathbf{x})|^{2}]\leq\sigma_{g}^{2}$, and either of the following conditions hold: ($f_{i}$ is monotonically non-decreasing and $\rho_{f}$-weakly convex, $g_{i}(\mathbf{x})$ is $\rho_{g}$-weakly convex; (ii)$f_{i}$ is $L_{\nabla f}$-smooth, $g_{i}(\mathbf{x})$ is $L_{\nabla g}$-smooth.
To tackle this objective, we also need to maintain and update estimators for inner functions $g_{i}(\mathbf{x})$. At the $t$-th iteration, we randomly sample an outer mini-batch $\mathcal{B}^{t}\in\{1,\ldots,n\}$, and draw inner mini-batch samples $\mathcal{B}_{1,i}^{t}$ for each $i\in\mathcal{B}^{t}$ to construct unbiased estimations ${g_{i}}(\mathbf{x}_{t};\mathcal{B}_{1,i}^{t})$. The variance-reduced estimator of $g_{i}(\mathbf{x}_{t})$ denoted by $u_{1,i}^{t}$ is updated by:
$u_{1,i}^{t+1}=(1-\gamma_{1})u_{1,i}^{t}+\gamma_{1}{g_{i}}(\mathbf{x}_{t};\mathcal{B}_{1,i}^{t})+\gamma_{1}^{\prime}({g_{i}}(\mathbf{x}_{t};\mathcal{B}_{1,i}^{t})-{g_{i}}(\mathbf{x}_{t-1};\mathcal{B}_{1,i}^{t})),$ (15)
where $\gamma_{1}\in(0,1),\gamma_{1}^{\prime}=\frac{n-|\mathcal{B}|}{|\mathcal{B}|(1-\gamma_{1})}+1-\gamma_{1}$, where $|\mathcal{B}|=|\mathcal{B}^{t}|$. Then, we approximate the gradient of $F(\mathbf{x}_{t})$ by:
$G_{1}^{t}=\frac{1}{|\mathcal{B}^{t}|}\sum_{i\in\mathcal{B}^{t}}\partial g_{i}(\mathbf{x}_{t};\mathcal{B}_{1,i}^{t})\partial f_{i}(u_{1,i}^{t}).$ (16)
Then we update $\mathbf{x}_{t+1}$ similarly as before by $\mathbf{x}_{t+1}=\mathbf{x}_{t}-\eta(G_{1}^{t}+G_{2}^{t})$. Due to the limit of space, we present full steps in Algorithm 2. The convergence of Algorithm 2 under condition (i) of Assumption 5.4 is stated below, and that under condition (ii) of Assumption 5.4 in the Appendix D.
Theorem 5.5.
Suppose Assumption 3.1, 5.1, 5.4 with condition (i) hold. Let $\gamma_{1}=\gamma_{2}=\mathcal{O}(\min\{|\mathcal{B}_{1,i}|,\frac{|\mathcal{B}_{2,k}|}{\beta^{2}}\}\frac{\epsilon^{4}}{\beta^{2}})$, $\eta=\mathcal{O}(\min\{\frac{|\mathcal{B}|}{n},\frac{|\mathcal{B}_{c}|}{\beta m}\}\min\{|\mathcal{B}_{1,i}|^{1/2},\frac{|\mathcal{B}_{2,k}|^{1/2}}{\beta}\}\frac{\epsilon^{4}}{\beta^{3}})$. After $T=\mathcal{O}(\max\{\frac{\beta}{|\mathcal{B}_{1,i}|^{1/2}},\frac{\beta^{2}}{|\mathcal{B}_{2,k}|^{1/2}},\frac{1}{|\mathcal{B}_{1,i}|}\}\max\{\frac{n}{|\mathcal{B}|},\frac{\beta m}{|\mathcal{B}_{c}|}\}\frac{\beta^{3}}{\epsilon^{6}})$ iterations, the Algorithm 2 satisfies $\mathbb{E}[\|\nabla\Phi_{\theta}(\mathbf{x}_{\hat{t}})\|]\leq\epsilon$ from some $\theta=O(1/(\rho_{0}+\rho_{1}\beta))$, where $\hat{t}$ is selected uniformly at random from $\{1,\ldots,T\}$
Remark: Combining the above result and that in Theorem 4.2, Algorithm 2 needs $T=\mathcal{O}(\frac{1}{\epsilon^{6}})$ to achieve a nearly $\epsilon$-KKT solution to the original problem (1). It is better than the complexity of $\mathcal{O}(1/\epsilon^{7})$ of the single-loop algorithm considered in [21].
6 Applications in Machine Learning
6.1 AUC Maximization with ROC Fairness Constraints
Figure 1: Training curves of 15 constraint functions of different methods for fair learning on the Adult dataset. Top: hinge-based penalty method with different $\beta$; Bottom: squared-hinge-based penalty method with different $\beta$. The legend train_tpr_th_-3.0 denotes the constraint function $h_{\tau^{+}}(\mathbf{w})$ with $\tau=3$; similarly, train_fpr_th_-3.0 represents the constraint function $h_{\tau^{-}}(\mathbf{w})$ with $\tau=3$.
Figure 2: Training curves of objective and constraint violation of different methods under a parameter setting when they satisfy the constraints in the end. Dashed lines correspond to objective AUC values (y-axis on the right) and solid lines correspond to constraint violations (y-axis on the left).
We consider learning a model with ROC fairness constraints [35]. Suppose the data are divided into two demographic groups $\mathcal{D}_{p}=\{(\mathbf{a}_{i}^{p},b_{i}^{p})\}_{i=1}^{n_{p}}$ and $\mathcal{D}_{u}=\{(\mathbf{a}_{i}^{u},b_{i}^{u})\}_{i=1}^{n_{u}}$, where $\mathbf{a}$ denotes the input data and $b\in\{1,-1\}$ denotes the class label. A ROC fairness is to ensure the ROC curves for classification of the two groups are the same. Since the ROC curve is constructed with all possible thresholds, we follow [35] by using a set of thresholds $\Gamma=\{\tau_{1},\cdots,\tau_{m}\}$ to define the ROC fairness. For each threshold $\tau$, we impose a constraint that the false positive rate (FPR) and true positive rate (TPR) of the two groups are close, formulated as the following:
$h_{\tau}^{+}(\mathbf{w}):=\Big{|}\frac{1}{n_{p}^{+}}\sum_{i=1}^{n_{p}}\mathbb{I}\{b^{p}_{i}=1\}\sigma(s_{\mathbf{w}}(\mathbf{a}_{i}^{p})-\tau)\vspace*{-0.1in}-\frac{1}{n_{u}^{+}}\sum_{i=1}^{n_{u}}\mathbb{I}\{b^{u}_{i}=1\}\sigma(s_{\mathbf{w}}(\mathbf{a}_{i}^{u})-\tau)\Big{|}-\kappa\leq 0$
$h_{\tau}^{-}(\mathbf{w}):=\Big{|}\frac{1}{n_{p}^{-}}\sum_{i=1}^{n_{p}}\mathbb{I}\{b^{p}_{i}=-1\}\sigma(s_{\mathbf{w}}(\mathbf{a}_{i}^{p})-\tau)\vspace*{-0.1in}-\frac{1}{n_{u}^{-}}\sum_{i=1}^{n_{u}}\mathbb{I}\{b^{u}_{i}=-1\}\sigma(s_{\mathbf{w}}(\mathbf{a}_{i}^{u})-\tau)\Big{|}-\kappa\leq 0,$
where $s_{\mathbf{w}}(\cdot)$ denotes a parameterized model, $\sigma(z)$ is the sigmoid function, and $\kappa>0$ is a tolerance parameter. For the objective function, we use a pairwise AUC loss:
$F(\mathbf{w})=\frac{-1}{|\mathcal{D}_{+}||\mathcal{D}_{-}|}\sum_{\mathbf{x}_{i}\in\mathcal{D}_{+}}\sum_{\mathbf{x}_{j}\in\mathcal{D}_{-}}\sigma(s(\mathbf{w},\mathbf{x}_{i})-s(\mathbf{w},\mathbf{x}_{j})),$
where $\mathcal{D}_{+}$ and $\mathcal{D}_{-}$ denote the set of positive and negative examples regardless their groups, respectively.
Our proposed penalty method solves the following problem:
$\min_{\mathbf{w}}F(\mathbf{w})+\frac{1}{2|{\Gamma}|}\sum_{\tau\in\Gamma}\beta[h_{\tau}^{+}(\mathbf{w})]_{+}+\beta[h_{\tau}^{-}(\mathbf{w})]_{+}.$ (17)
It is notable that the above problem does not directly satisfy the Assumption 5.1. However, the Algorithm 1 with a minor change is still applicable by formulating $[h_{\tau}^{+}(\mathbf{w})]_{+}=f(g(h_{\tau,1}(\mathbf{w})))$, where $f(g)=[g]_{+}$, $g(h)=|h|-\kappa$ and $h_{\tau,1}(\mathbf{w})=\frac{1}{n_{p}^{+}}\sum_{i=1}^{n_{p}}\mathbb{I}\{b_{i}^{p}=1\}\sigma(s_{\mathbf{w}}(\mathbf{a}_{i}^{p})-\tau)$. As a result, $f$ is monotonically non-decreasing and convex, $g$ is convex, and $h_{\tau,1}$ is a smooth function, which fits another setting in [16]. The only change to Algorithm 1 is to maintain an estimator of $h_{\tau,1},\tau\in\Gamma$ and handle $[h_{\tau}^{-}(\mathbf{w})]_{+}$ similarly.
Setting. For experiments, we use three datasets, Adult and COMPAS used in [12], CheXpert [18], which contain male/female, Caucasian/non-Caucasian, and male/female groups, respectively. We set $\Gamma=\{-3,-2,-1,\cdots,3\}$. For models, we use a simple neural network with 2 hidden layers for Adult and COMPAS data and use DenseNet121 for CheXpert data [40]. We implement Algorithm 1 with setting $\gamma_{2}=0.8$, $\gamma_{2}^{\prime}=0.1$, $\kappa=0.005$. We compare with the algorithm in [21] which optimizes a squared-hinge penalty function. For both algorithms, we tune the initial learning rate in {1e-3, 1e-4} and decay it at 50% and 75% epochs by a factor of 10. We tuned $\beta=\{1,4,8,10,20,40,80,100\}$ for our hinge exact penalty method and $\beta=\{10,40,80,100,200,400,800,1000\}$ for squared hinge penalty method. We also compare a double-loop method (ICPPAC) [6, Algorithm 4] for Adult and COMPAS, where we tune their $\eta$ in {0.1, 0.01}, $\tau$ in {1, 10, 100}, $\mu$ in {1e-2, 1e-3, 1e-4}, and fix $\theta_{t}$ to 0.1. We do not report ICPPAC on the large-scale CheXpert data because it is not efficient due to back-propagation on 15 individual constraint functions every iteration. We run each method for total 60 epoches with a batch size of 128. As optimization involves randomness, we run all the experiments with five different random seeds and then calculate average of the AUC scores and constraint values.
Figure 3: Training curves of 5 constraint values in zero-one loss of different methods for continual learning with non-forgetting constraints when targeting the foggy class. Top: hinge penalty method with different $\beta$; Bottom: squared-hinge penalty method with different $\beta$.
Figure 4: Comparison of training curves between the hinge penalty method and the squared-hinge penalty method when both methods satisfy the constraints in the end. Dashed lines correspond to target accuracy improvement over the base model (y-axis on the right) and solid lines correspond to constraint violation in terms of zero-one loss (y-axis on the left).
Result. We present the training curves of values of constraint functions in Fig. 1 on Adult dataset to illustrate which $\beta$ value satisfies the constraints. For the hinge exact penalty method, when $\beta=20$, all constraints are satisfied. In contrast, the squared hinge penalty method requires a much larger $\beta=800$ to satisfy the constraints at the end. We include more results on COMPAS and CheXpert in Appendix A. We compare the training curves of objective AUC values and the constraint violation as measured by the worst constraint function value at each epoch of different methods in Fig. 2 on different datasets. These results demonstrate that the hinge exact penalty method has a better performance than the squared hinge penalty method in terms of the objective AUC value when both have a similar constraint satisfaction, and is competitive with if not better than ICPPAC method.
6.2 Continual learning with non-forgetting constraints
Continual learning with non-forgetting constraints has been considered in [21], which is termed the model developmental safety. We consider the same problem of developing the CLIP model while satisfying developmental safety constraints:
$\min_{\mathbf{w}}\ F(\mathbf{w},\mathcal{D}):=\frac{1}{n}\sum_{(\mathbf{x}_{i},\mathbf{t}_{i})\in\mathcal{D}}L_{\text{ctr}}(\mathbf{w},\mathbf{x}_{i},\mathbf{t}_{i},\mathcal{T}_{i}^{-},\mathcal{I}_{i}^{-}),$
[…]
The $L_{\text{ctr}}(\mathbf{w},\mathbf{x}_{i},\mathbf{t}_{i},\mathcal{T}_{i}^{-},\mathcal{I}_{i}^{-})$ is a two-way contrastive loss for each image-text pair $(\mathbf{x}_{i},\mathbf{t}_{i})$ [41],
$L_{\text{ctr}}(\mathbf{w};\mathbf{x}_{i},t_{i},\mathcal{T}_{i}^{-},\mathcal{I}_{i}^{-}):$ $=-\tau\log\frac{\exp(E_{1}(\mathbf{w},\mathbf{x}_{i})^{\top}E_{2}(\mathbf{w},t_{i})/\tau)}{\sum_{t_{j}\in\mathcal{T}_{i}^{-}}\exp(E_{1}(\mathbf{w},\mathbf{x}_{i})^{\top}E_{2}(\mathbf{w},t_{j})/\tau)}$
$~{}~{}~{}~{}~{}~{}-\tau\log\frac{\exp(E_{2}(\mathbf{w},t_{i})^{\top}E_{1}(\mathbf{w},\mathbf{x}_{i})/\tau)}{\sum_{\mathbf{x}_{j}\in\mathcal{I}_{i}^{-}}\exp(E_{2}(\mathbf{w},t_{j})^{\top}E_{1}(\mathbf{w},\mathbf{x}_{i})/\tau)},$
where $E_{1}(\mathbf{w},\mathbf{x})$ and $E_{2}(\mathbf{w},t)$ denotes a (normalized) encoded representation of an image $\mathbf{x}$ and a text $t$, respectively. $\mathcal{T}_{i}^{-}$ denotes the set of all texts to be contrasted with respect to (w.r.t) $\mathbf{x}_{i}$ (including itself) and $\mathcal{I}_{i}^{-}$ denotes the set of all images to be contrasted w.r.t $t_{i}$ (including itself). Here, the data $\mathcal{D}$ is a target dataset. $\mathcal{L}_{k}(\mathbf{w},\mathcal{D}_{k})=\frac{1}{n_{k}}\sum_{(\mathbf{x}_{i},y_{i})\sim\mathcal{D}_{k}}\ell_{k}(\mathbf{w},\mathbf{x}_{i},y_{i})$, where $\ell_{k}$ is a logistic loss for the $k$-th classification task. We use our Algorithm 2 to optimize the hinge penalty function.
Setting. We follow the exactly same experiment setting as in [21]. We experiment on the large-scale diverse driving image dataset, namely BDD100K [38]. This dataset involves classification of six weather conditions, i.e., clear, overcast, snowy, rainy, partly cloudy, foggy, and of six scene types, i.e., highway, residential area, city street, parking lot, gas station, tunnel. We consider three objectives of contrastive loss for finetuning a pretrained CLIP model to target at improving classification of foggy, tunnel and overcast, respectively. For each task, we use other weather or scene types to formulate the constraints, ensuring the new model does not lose the performance on these other classes. The data of the objective for the target class are sampled from the training set of BDD100K and the external LAION400M dataset [32] and the data of constrained tasks are sample from BDD100K, where the statistics are given in [21].
For all methods, we fix the learning rate $1e^{-6}$ with cosine scheduler, using a weight decay of 0.1. We run each method for a total of 60 epochs with a batch size of 256 and 400 iterations per epoch. We tune $\beta$ in $\{0.01,0.1,1,5,10,20\}$ for our method and in $\{80,100,200,400,800\}$ for squared hinge penalty method and set $|\mathcal{B}_{c}|=m,|\mathcal{B}_{k}|=10,\gamma_{1}=\gamma_{2}=0.8,\gamma_{1}^{\prime}=\gamma_{2}^{\prime}=0.1,\tau=0.05$.
Results. We present the training curves of the constraint functions as measured by zero-one loss in Fig. 4 for different methods with different $\beta$ values on targeting foggy class. We can see that when $\beta=10$, our hinge penalty method will satisfy all constraints. In contrast, the squared hinge penalty method needs a much larger $\beta=800$ to satisfy all constraints. The constraint curves when targeting overcast and tunnel are presented in Figs. 5, 6 in Appendix A.
Fig. 4 compares the training curves of the target accuracy improvement between our hinge penalty method and squared-hinge penalty method when both satisfy the constraints. According to the Fig. 4, our hinge penalty method converges faster in terms of the objective measure when they exhibit similar constraint satisfaction in the end.
7 Conclusion
In this paper, we have studied non-convex constrained optimization with a weakly-convex objective and weakly convex constraint functions, which are all stochastic. We developed a hinge-based exact penalty method and its theory to guarantee finding a nearly $\epsilon$-KKT solution. By leveraging stochastic optimization techniques for non-smooth weakly-convex finite-sum coupled compositional optimization problems, we developed algorithms for solving the penalty function with different structures of the objective functions. Our experiments on fair learning with ROC fairness constraints and continual learning with non-forgetting constraints demonstrate the effectiveness of our algorithms compared.
Appendix A More Experiment Results
The experiments of AUC Maximization with ROC Fairness Constraints in our paper is run on a GPU server with 10 A30 24G GPUs. The experiment of continual learning with non-forgetting constraints is run on High Performance Research Computing platforms, which contains 117 GPU nodes, each with two A100 40GB GPUs.
We give other experiment results in this Appendix. First, we compute the minimum singular value $\delta$ of the Jacobian matrix of violating constraint functions in our experiments at different solutions to verify the conditions in Lemma 4.4, as shown in the Table 2. For the first experiment, we compute $\sigma$ based on the Adult dataset at the initial model, final model, and two randomly selected intermediate epochs. For our second experiment, we compute $\sigma$ when Target Foggy at both the initial and final models. All results demonstrate that the minimum singular value $\sigma$ of $H(\mathbf{x})$ remains positive, which is consistent with the theoretical guarantee stated in Lemma 4.4.
Experiment 1 (Adult) Initial Model Epoch_30 Epoch_45 Final Model
Minimum Singular Values $\sigma$ of $\nabla H(\mathbf{x})$ 0.000648 0.000112 0.000102 0.000101
Experiment 2 (Target foggy) Initial Model Final Model
Minimum Singular Values $\delta$ of $H(\mathbf{x})$ 24.1038 16.3397
Table 2: Minimum singular values $\delta$ of the Jacobian matrix $H(\mathbf{x})$ at different solution.
For targeting overcast we consider other weather conditions except for foggy as protected tasks due to that there is a lack of foggy data in BDD100k for defining a significant constraint. For the same reason, we consider other scence types except gas station as protected tasks for targeting tunnel. We use the mostly same experiment setting with the case target foggy only difference when tuning $\beta$. We tune $\beta=\{0.01,0.1,1,10,20\}$ for hinge-based penalty method and $\beta=\{0.1,1,20,50,100,200\}$ for squared hinge penalty method when target overcast. When target tunnel, we tune $\beta=\{0.01,0.1,1,10,20\}$ for both hinge-based penalty method for squared hinge penalty method. The training curves of target overcast and tunnel show in Fig. 5 and Fig. 6. The training curves on dataset COMPAS and Chexpert show in Fig. 7 and Fig. 8 for experiment ROC Fairness Constraints.
Figure 5: Training curves of 5 constraint values in zero-one loss of different methods for continual learning with non-forgetting constraints when targeting the overcast class. Top: hinge penalty method with different $\beta$; Bottom: squared-hinge penalty method with different $\beta$.
Figure 6: Training curves of 5 constraint values in zero-one loss of different methods for continual learning with non-forgetting constraints when targeting the tunnel class. Top: hinge penalty method with different $\beta$; Bottom: squared-hinge penalty method with different $\beta$.
Figure 7: Training curves of 15 constraint functions of different methods for fair learning on the COMPAS dataset. Top: hinge-based penalty method with different $\beta$; Bottom: squared-hinge-based penalty method with different $\beta$.
Figure 8: Training curves of 15 constraint functions of different methods for fair learning on the CheXpert dataset. Top: hinge-based penalty method with different $\beta$; Bottom: squared-hinge-based penalty method with different $\beta$.
Appendix B Technique Lemmas
To prove our main theorem, we need the following lemmas.
Lemma B.1.
[Lemma 1 [20]] Consider update (15). Under Assumptions 3.1, 5.1 and 5.4, by setting $\gamma_{1}^{\prime}=\frac{n-|\mathcal{B}|}{|\mathcal{B}|(1-\gamma_{1})}+(1-\gamma_{1})$, for $\gamma_{1}\leq\frac{1}{2}$, the function value variance $\Xi^{t+1}:=\frac{1}{n}\sum_{i=1}^{n}\|u_{1,i}^{t+1}-g_{i}(\mathbf{x}_{t+1})\|^{2}$ can be bounded as
$\mathbb{E}\left[\Xi^{t+1}\right]\leq(1-\frac{|\mathcal{B}|\gamma_{1}}{n})\mathbb{E}\left[\Xi^{t}\right]+\frac{8nL_{g}^{2}}{|\mathcal{B}|}\mathbb{E}[\|\mathbf{x_{t+1}}-\mathbf{x_{t}}\|^{2}]+\frac{2\gamma_{1}^{2}|\mathcal{B}|\sigma_{g}^{2}}{n|\mathcal{B}_{1,i}|}.$ (18)
Lemma B.2.
[Lemma 1 [20]] Consider update (12). Under Assumptions 3.1 and 5.1, by setting $\gamma_{2}^{\prime}=\frac{m-|\mathcal{B}_{c}|}{|\mathcal{B}_{c}|(1-\gamma_{2})}+(1-\gamma_{2})$, for $\gamma_{2}\leq\frac{1}{2}$, the function value variance $\Gamma^{t+1}:=\frac{1}{m}\sum_{k=1}^{m}\|u_{2,k}^{t+1}-h_{k}(\mathbf{x}_{t+1})\|^{2}$ can be bounded as
$\mathbb{E}\left[\Gamma_{t+1}\right]\leq(1-\frac{|\mathcal{B}_{c}|\gamma_{2}}{m})\mathbb{E}[\Gamma_{t}]+\frac{8mL_{h}^{2}}{|\mathcal{B}_{c}|}\mathbb{E}[\|\mathbf{x_{t+1}}-\mathbf{x_{t}}\|^{2}]+\frac{2\gamma_{1}^{2}|\mathcal{B}_{c}|\sigma_{h}^{2}}{m|\mathcal{B}_{2,k}|}.$ (19)
From the analysis of Lemma 4.5 in [16], the following two inequalities hold true based on Lemma B.1 and Lemma B.2. For simplicity, we use a constant $M^{2}\geq\max\{2L_{f}^{2}L_{g}^{2}+2\beta^{2}L_{h}^{2},2\sigma_{f}^{2}+2L_{F}^{2}+2\beta^{2}L_{h}^{2}\}$.
$\mathbb{E}[\frac{1}{m}\sum_{k=1}^{m}\|{h_{k}}(\mathbf{x}_{t})-u_{2,k}^{t}\|]\leq(1-\frac{|\mathcal{B}_{c}|\gamma_{2}}{2m})^{t}\frac{1}{m}\sum_{k=1}^{m}\|u_{2,k}^{0}-h_{k}(\mathbf{x}_{0})\|+\frac{4mL_{h}\eta M}{|\mathcal{B}_{c}|\gamma_{2}^{1/2}}+\frac{2\gamma_{2}^{1/2}\sigma_{h}}{|\mathcal{B}_{2,k}|^{1/2}},$ (20)
and
$\mathbb{E}[\frac{1}{n}\sum_{i=1}^{n}\|g_{i}(\mathbf{x}_{t})-u_{1,i}^{t}\|]\leq(1-\frac{|\mathcal{B}|\gamma_{1}}{2n})^{t}\frac{1}{n}\sum_{i=1}^{n}\|u_{1,i}^{0}-g_{i}(\mathbf{x}_{0})\|+\frac{4nL_{g}M\eta}{|\mathcal{B}|\gamma_{1}^{1/2}}+\frac{2\gamma_{1}^{1/2}\sigma_{g}}{|\mathcal{B}_{1,i}|^{1/2}}.$ (21)
Proof of Lemma 4.1.
Notice that we have assumed $F(\cdot)$ to be $\rho_{0}$-weakly convex in Assumption 3.1. By the definition of weak convexity, it suffices to show $[h_{k}(\mathbf{x})]_{+}$ to be $\rho_{1}$-weakly convex for all $k=1,\dots,m$.
Let $[h_{k}(x)]^{\prime}_{+}$ denote any element in $\partial[h_{k}(x)]_{+}$. Given any $k\in\{1,\dots,m\}$ and $x,x^{\prime}$, following from the convexity of $[\cdot]_{+}$, we have
${}_{+}-[h_{k}(x^{\prime})]_{+}$ $\geq\partial[h_{k}(x^{\prime})]_{+}(h(x)-h(x^{\prime}))$ (22)
$\stackrel{{\scriptstyle(a)}}{{\geq}}\partial[h_{k}(x^{\prime})]_{+}(\langle\partial h_{k}(x^{\prime}),x-x^{\prime}\rangle-\frac{\rho_{1}}{2}\|x-x^{\prime}\|^{2})$
$\stackrel{{\scriptstyle(b)}}{{\geq}}\langle\partial h_{k}(x^{\prime})\partial[h_{k}(x^{\prime})]_{+},x-x^{\prime}\rangle-\frac{\rho_{1}}{2}\|x-x^{\prime}\|^{2},$
where $(a)$ follows from the fact $[h_{k}(x^{\prime})]^{\prime}_{+}\geq 0$ for all $x^{\prime}$ and the $\rho_{1}$-weak convexity of $h_{k}(\cdot)$, and $(b)$ follows from the fact $[h_{k}(x^{\prime})]^{\prime}_{+}\leq 1$.
To show the $L$-Lipschitz continuity of $\Phi(\cdot)$, we utilize the $L_{F}$ and $L_{h}$-Lipschitz continuity of $F(\cdot)$ and $h_{k}(\cdot)$ to obtain that for any $x,x^{\prime}$ we have
$\|\Phi(x)-\Phi(x^{\prime})\|$ $\leq\|F(x)-F(x^{\prime})\|+\frac{\beta}{m}\sum_{k=1}^{m}\|[h_{k}(x)]_{+}-[h_{k}(x^{\prime})]_{+}\|$ (23)
$\leq L_{F}\|x-x^{\prime}\|+\frac{\beta}{m}\sum_{k=1}^{m}\|h_{k}(x)-h_{k}(x^{\prime})\|$
$\leq L_{F}\|x-x^{\prime}\|+\frac{\beta}{m}\sum_{k=1}^{m}L_{h}\|x-x^{\prime}\|$
$\leq(L_{F}+\beta L_{h})\|x-x^{\prime}\|.$
Proof of Lemma 4.3.
Our goal is to prove that there exists a constant $\delta>0$ for any $\mathbf{x}$ such that $h(\mathbf{x})>0$ (violating the constraint), we have $\|\nabla h(\mathbf{x})\|\geq\delta$.
By conditions (i) and (ii) and the fact that $h(x)>0$, we have
$c<h(\mathbf{x})-\min_{\mathbf{y}}h(\mathbf{y})\leq\frac{\|\nabla h(\mathbf{x})\|^{2}}{2\mu}.$
This means $\|\nabla h(\mathbf{x})\|\geq\sqrt{2c\mu}=:\delta$. ∎
Proof of Lemma 4.4.
Let $\boldsymbol{\xi}=(\xi_{1},\ldots,\xi_{m})^{T}$, where $\xi_{k}$ defined as
$\xi_{k}=\begin{cases}1&\text{if }h_{k}({\mathbf{x}})>0,\\ [0,1]&\text{if }h_{k}({\mathbf{x}})=0,\\ 0&\text{if }h_{k}({\mathbf{x}})<0,\end{cases}\quad\in[h_{k}({\mathbf{x}})]^{\prime}_{+}.$
Then
$\left\{\sum_{k=1}^{m}\xi_{k}\nabla h_{k}(\mathbf{x})\right\}=\left\{\nabla H(\mathbf{x})^{\top}\boldsymbol{\xi}\right\}\subseteq\sum_{k=1}^{m}\partial[h_{k}(\mathbf{\mathbf{x}})]_{+},$
where $\nabla H(\mathbf{x})=(\nabla h_{1}(\mathbf{x}),\ldots,\nabla h_{k}(\mathbf{x}))^{T}.$ Then for $\forall\mathbf{x}\in\mathcal{V}$, where $\mathcal{V}=\{\mathbf{x}:\max_{k}h_{k}(\mathbf{x})>0\},$
$\operatorname{dist}\left(0,\frac{1}{m}\sum_{k=1}^{m}\partial[h_{k}(\mathbf{x})]_{+}\right)\geq\left\|\frac{1}{m}\nabla H(\mathbf{x})^{\top}\boldsymbol{\xi}\right\|\geq\frac{1}{m}\lambda_{\min}(\nabla H(\mathbf{x}))\|\boldsymbol{\xi}\|\geq\frac{\sigma}{m}.$
Appendix C Proof of Theorem 5.3
Proof.
The objective function
$\min_{\mathbf{x}}\Phi(\mathbf{x}):=\underbrace{\mathbb{E}_{\zeta}[f(\mathbf{x};\zeta)]}_{F(\mathbf{x})}+\underbrace{\frac{\beta}{m}\sum_{k=1}^{m}[h_{k}(\mathbf{x})]_{+}}_{H(\mathbf{x})}.$ (24)
For simplicity, denote $\bar{\mathbf{x}}_{t}:=\text{prox}_{\Phi/\bar{\rho}}(\mathbf{x}_{t})$. Consider change in the Moreau Envelope
$\mathbb{E}_{t}[\Phi_{1/\bar{\rho}}(\mathbf{x}_{t+1})]$ $=\mathbb{E}_{t}[\min_{\tilde{\mathbf{x}}}\Phi(\tilde{\mathbf{x}})+\frac{\bar{\rho}}{2}\|\tilde{\mathbf{x}}-\mathbf{x}_{t+1}\|^{2}]$
$\leq\mathbb{E}_{t}[\Phi(\bar{\mathbf{x}}_{t})+\frac{\bar{\rho}}{2}\|\bar{\mathbf{x}}_{t}-\mathbf{x}_{t+1}\|^{2}]$
$=\mathbb{E}_{t}[\Phi(\bar{\mathbf{x}}_{t})+\frac{\bar{\rho}}{2}\|\bar{\mathbf{x}}_{t}-(\mathbf{x}_{t}-\eta(G_{1}^{t}+G_{2}^{t}))\|^{2}]$
$\leq\Phi(\bar{\mathbf{x}}_{t})+\frac{\bar{\rho}}{2}\|\bar{\mathbf{x}}_{t}-\mathbf{x}_{t}\|^{2}+\bar{\rho}\mathbb{E}_{t}[\eta\langle\bar{\mathbf{x}}_{t}-\mathbf{x}_{t},G_{1}^{t}+G_{2}^{t}\rangle]+\eta^{2}\frac{\bar{\rho}M^{2}}{2}$
$=\Phi_{1/\bar{\rho}}(\mathbf{x}_{t})+\bar{\rho}\eta\langle\bar{\mathbf{x}}_{t}-\mathbf{x}_{t},\mathbb{E}_{t}[G_{1}^{t}]\rangle+\bar{\rho}\eta\langle\bar{\mathbf{x}}_{t}-\mathbf{x}_{t},\mathbb{E}_{t}[G_{2}^{t}]\rangle+\eta^{2}\frac{\bar{\rho}M^{2}}{2},$ (25)
where the second inequality uses the bound of $\mathbb{E}_{t}[\|G_{1}^{t}+G_{2}^{t}\|^{2}]$, which follows from the Lipchitz continuity and bounded variance, denoted by $M^{2}\geq 2\sigma_{f}^{2}+2L_{F}^{2}+2\beta^{2}L_{h}^{2}$. Here,
$\mathbb{E}_{t}[G_{1}^{t}]\in\partial F(\mathbf{x}_{t}),~{}~{}~{}~{}\mathbb{E}_{t}[G_{2}^{t}]\in\frac{\beta}{m}\sum_{k=1}^{m}\partial h_{k}(\mathbf{x}_{t})[u_{2,k}^{t}]^{\prime}_{+}.$
Next we give the bound of $\langle\bar{\mathbf{x}}_{t}-\mathbf{x}_{t},\mathbb{E}_{t}[G_{1}^{t}]\rangle$ and $\langle\bar{\mathbf{x}}_{t}-\mathbf{x}_{t},\mathbb{E}_{t}[G_{2}^{t}]\rangle$. To this end, first we give the bound of $\langle\bar{\mathbf{x}}_{t}-\mathbf{x}_{t},\mathbb{E}_{t}[G_{1}^{t}]\rangle$. Since $F$ is $\rho_{0}$-weakly convex, we have
$F(\bar{\mathbf{x}}_{t})-F(\mathbf{x}_{t})\geq\partial F(\mathbf{x}_{t})^{\top}(\bar{\mathbf{x}}_{t}-\mathbf{x}_{t})-\frac{\rho_{0}}{2}\|\bar{\mathbf{x}}_{t}-\mathbf{x}_{t}\|^{2}.$
Then it follows
$\partial F(\mathbf{x}_{t})^{\top}(\bar{\mathbf{x}}_{t}-\mathbf{x}_{t})\leq F(\bar{\mathbf{x}}_{t})-F(\mathbf{x}_{t})+\frac{\rho_{0}}{2}\|\bar{\mathbf{x}}_{t}-\mathbf{x}_{t}\|^{2}.$ (26)
Next we bound $\langle\bar{\mathbf{x}}_{t}-\mathbf{x}_{t},\mathbb{E}_{t}[G_{2}^{t}]\rangle$. For given $k\in\{1,\ldots,m\}$, we get
$[{h_{k}}(\bar{\mathbf{x}}_{t})]_{+}-[u_{2,k}^{t}]_{+}\geq[u_{2,k}^{t}]^{\prime}_{+}({h_{k}}(\bar{\mathbf{x}}_{t})-u_{2,k}^{t})$
$\geq[u_{2,k}^{t}]^{\prime}_{+}\left[{h_{k}}(\mathbf{x}_{t})-u_{2,k}^{t}+\partial{h_{k}}(\mathbf{x}_{t})^{\top}(\bar{\mathbf{x}}_{t}-\mathbf{x}_{t})-\frac{\rho_{1}}{2}\|\bar{\mathbf{x}}_{t}-\mathbf{x}_{t}\|^{2}\right]$
$\geq[u_{2,k}^{t}]^{\prime}_{+}({h_{k}}(\mathbf{x}_{t})-u_{2,k}^{t})+[u_{2,k}^{t}]^{\prime}_{+}\partial{h_{k}}(\mathbf{x}_{t})^{\top}(\bar{\mathbf{x}}_{t}-\mathbf{x}_{t})-\frac{\rho_{1}}{2}\|\bar{\mathbf{x}}_{t}-\mathbf{x}_{t}\|^{2},$
where the first inequality uses the convexity of function $[\cdot]_{+}$, the second inequality uses the fact $[u_{2,k}^{t}]^{\prime}_{+}\geq 0$ and the $\rho_{1}$-weak convexity of $h_{k}(\cdot)$, and the last inequality uses the fact that $0\leq[u_{2,k}^{t}]^{\prime}_{+}\leq 1$. Then it follows
$\beta\frac{1}{m}\sum_{k=1}^{m}[u_{2,k}^{t}]^{\prime}_{+}\partial{h_{k}}(\mathbf{x}_{t})^{\top}(\bar{\mathbf{x}}_{t}-\mathbf{x}_{t})$
$\leq\frac{\beta}{m}\sum_{k=1}^{m}\Big{[}[{h_{k}}(\bar{\mathbf{x}}_{t})]_{+}-[u_{2,k}^{t}]_{+}-[u_{2,k}^{t}]^{\prime}_{+}({h_{k}}(\mathbf{x}_{t})-u_{2,k}^{t})+\frac{\rho_{1}}{2}\|\bar{\mathbf{x}}_{t}-\mathbf{x}_{t}\|^{2}\Big{]}.$ (27)
Adding above two estimation (26) and (27) back to (25), we have
$\mathbb{E}_{t}[\Phi_{1/\bar{\rho}}(\mathbf{x}_{t+1})]$ $\leq\Phi_{1/\bar{\rho}}(\mathbf{x}_{t})+\bar{\rho}\eta\langle\bar{\mathbf{x}}_{t}-\mathbf{x}_{t},\mathbb{E}_{t}[G_{1}^{t}]\rangle+\bar{\rho}\eta\langle\bar{\mathbf{x}}_{t}-\mathbf{x}_{t},\mathbb{E}_{t}[G_{2}^{t}]\rangle+\eta^{2}\frac{\bar{\rho}M^{2}}{2}$
$\leq\Phi_{1/\bar{\rho}}(\mathbf{x}_{t})+\eta^{2}\frac{\bar{\rho}M^{2}}{2}+\bar{\rho}\eta\Big{[}F(\bar{\mathbf{x}}_{t})-F(\mathbf{x}_{t})+\frac{\rho_{0}}{2}\|\bar{\mathbf{x}}_{t}-\mathbf{x}_{t}\|^{2}\Big{]}$
$+\bar{\rho}\eta\frac{\beta}{m}\sum_{k=1}^{m}\Big{[}[{h_{k}}(\bar{\mathbf{x}}_{t})]_{+}-[u_{2,k}^{t}]_{+}-[u_{2,k}^{t}]^{\prime}_{+}({h_{k}}(\mathbf{x}_{t})-u_{2,k}^{t})+\frac{\rho_{1}}{2}\|\bar{\mathbf{x}}_{t}-\mathbf{x}_{t}\|^{2}\Big{]}$
$=\Phi_{1/\bar{\rho}}(\mathbf{x}_{t})+\eta^{2}\frac{\bar{\rho}M^{2}}{2}+\bar{\rho}\eta\Big{[}F(\bar{\mathbf{x}}_{t})-F(\mathbf{x}_{t})+\frac{\rho_{0}}{2}\|\bar{\mathbf{x}}_{t}-\mathbf{x}_{t}\|^{2}\Big{]}$
$+\bar{\rho}\eta\frac{\beta}{m}\sum_{k=1}^{m}\Big{[}[{h_{k}}(\bar{\mathbf{x}}_{t})]_{+}-[{h_{k}}(\mathbf{x}_{t})]_{+}+[{h_{k}}(\mathbf{x}_{t})]_{+}-[u_{2,k}^{t}]_{+}-[u_{2,k}^{t}]^{\prime}_{+}({h_{k}}(\mathbf{x}_{t})-u_{2,k}^{t})+\frac{\rho_{1}}{2}\|\bar{\mathbf{x}}_{t}-\mathbf{x}_{t}\|^{2}\Big{]}.$ (28)
By Lemma 4.1, the function $\Phi$ is $C$-weakly convex. We have $(\bar{\rho}-C)-$strong convexity of $\mathbf{x}\mapsto\Phi(\mathbf{x})+\frac{\bar{\rho}}{2}\|\mathbf{x}_{t}-\mathbf{x}\|^{2}$
$\Phi(\mathbf{\bar{x}}_{t})-\Phi(\mathbf{x}_{t})$ $=\Big{(}\Phi(\mathbf{\bar{x}}_{t})+\frac{\bar{\rho}}{2}\|\mathbf{x}_{t}-\bar{\mathbf{x}}_{t}\|^{2}\Big{)}-\Big{(}\Phi(\mathbf{{x}}_{t})+\frac{\bar{\rho}}{2}\|\mathbf{x}_{t}-{\mathbf{x}}_{t}\|^{2}\Big{)}-\frac{\bar{\rho}}{2}\|\mathbf{\bar{x}}_{t}-\mathbf{x}_{t}\|^{2}$
$\leq(\frac{C}{2}-\bar{\rho})\|\mathbf{\bar{x}}_{t}-\mathbf{x}_{t}\|^{2}.$
Then we get
$F(\bar{\mathbf{x}}_{t})-F(\mathbf{x}_{t})+\beta\frac{1}{m}\sum_{k=1}^{m}([{h_{k}}(\bar{\mathbf{x}}_{t})]_{+}-[{h_{k}}(\mathbf{x}_{t})]_{+})=\Phi(\mathbf{\bar{x}}_{t})-\Phi(\mathbf{x}_{t})\leq(\frac{C}{2}-\bar{\rho})\|\mathbf{\bar{x}}_{t}-\mathbf{x}_{t}\|^{2}.$ (29)
Plugging inequality (29) to (28), we get
$\mathbb{E}_{t}[\Phi_{1/\bar{\rho}}(\mathbf{x}_{t+1})]$
$\leq\Phi_{1/\bar{\rho}}(\mathbf{x}_{t})+\eta^{2}\frac{\bar{\rho}M^{2}}{2}+\bar{\rho}\eta\Big{(}\frac{\rho_{0}}{2}+\beta\frac{\rho_{1}}{2}+\frac{C}{2}-\bar{\rho}\Big{)}\|\bar{\mathbf{x}}_{t}-\mathbf{x}_{t}\|^{2}$
$+\bar{\rho}\eta\frac{\beta}{m}\sum_{k=1}^{m}\Big{[}[{h_{k}}(\mathbf{x}_{t})]_{+}-[u_{2,k}^{t}]_{+}-[u_{2,k}^{t}]^{\prime}_{+}({h_{k}}(\mathbf{x}_{t})-u_{2,k}^{t})\Big{]}.$
Setting $\frac{\rho_{0}}{2}+\beta\frac{\rho_{1}}{2}+\frac{C}{2}=\frac{\bar{\rho}}{2}$, we have
$\mathbb{E}_{t}[\Phi_{1/\bar{\rho}}(\mathbf{x}_{t+1})]$
$\leq\Phi_{1/\bar{\rho}}(\mathbf{x}_{t})+\eta^{2}\frac{\bar{\rho}M^{2}}{2}-\eta\frac{\bar{\rho}^{2}}{2}\|\bar{\mathbf{x}}_{t}-\mathbf{x}_{t}\|^{2}+\bar{\rho}\eta\frac{\beta}{m}\sum_{k=1}^{m}\Big{[}[{h_{k}}(\mathbf{x}_{t})]_{+}-[u_{2,k}^{t}]_{+}-[u_{2,k}^{t}]^{\prime}_{+}({h_{k}}(\mathbf{x}_{t})-u_{2,k}^{t})\Big{]}$
$\leq\Phi_{1/\bar{\rho}}(\mathbf{x}_{t})+\eta^{2}\frac{\bar{\rho}M^{2}}{2}-\frac{\eta}{2}\|\nabla\Phi_{1/\bar{\rho}}(\mathbf{x}_{t})\|^{2}+\bar{\rho}\eta\frac{\beta}{m}\sum_{k=1}^{m}\Big{[}[{h_{k}}(\mathbf{x}_{t})]_{+}-[u_{2,k}^{t}]_{+}-[u_{2,k}^{t}]^{\prime}_{+}({h_{k}}(\mathbf{x}_{t})-u_{2,k}^{t})\Big{]}.$
Since $0\leq[u_{2,k}^{t}]^{\prime}_{+}\leq 1$, we get
$\mathbb{E}_{t}[\Phi_{1/\bar{\rho}}(\mathbf{x}_{t+1})]\leq\Phi_{1/\bar{\rho}}(\mathbf{x}_{t})+\eta^{2}\frac{\bar{\rho}M^{2}}{2}-\frac{\eta}{2}\|\nabla\Phi_{1/\bar{\rho}}(\mathbf{x}_{t})\|^{2}+2\bar{\rho}\eta\frac{\beta}{m}\sum_{k=1}^{m}\|{h_{k}}(\mathbf{x}_{t})-u_{2,k}^{t}\|.$ (30)
Taking the full expectation on both sides of (30) and applying the inequality of (20), we have
$\mathbb{E}[\Phi_{1/\bar{\rho}}(\mathbf{x}_{t+1})]$
$\leq\mathbb{E}[\Phi_{1/\bar{\rho}}(\mathbf{x}_{t})]+\eta^{2}\frac{\bar{\rho}M^{2}}{2}-\frac{\eta}{2}\mathbb{E}[\|\nabla\Phi_{1/\bar{\rho}}(\mathbf{x}_{t})\|^{2}]$
$+2\bar{\rho}\eta\beta\Big{[}(1-\frac{|\mathcal{B}_{c}|\gamma_{2}}{2m})^{t}\frac{1}{m}\sum_{k=1}^{m}\|u_{2,k}^{0}-h_{k}(\mathbf{x}_{0})\|+\frac{4mL_{h}\eta M}{|\mathcal{B}_{c}|\gamma_{2}^{1/2}}+\frac{2\gamma_{2}^{1/2}\sigma_{h}}{|\mathcal{B}_{2,k}|^{1/2}}\Big{]}.$
Taking summation from $t=0$ to $T-1$ yields
$\mathbb{E}[\Phi_{1/\bar{\rho}}(\mathbf{x}_{T})]$ $\leq\Phi_{1/\bar{\rho}}(\mathbf{x}_{0})+\eta^{2}T\frac{\bar{\rho}M^{2}}{2}-\frac{\eta}{2}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla\Phi_{1/\bar{\rho}}(\mathbf{x}_{t})\|^{2}]$
$+2\bar{\rho}\eta\beta\Big{[}\frac{2m}{|\mathcal{B}_{c}|\gamma_{2}}\frac{1}{m}\sum_{k=1}^{m}\|u_{2,k}^{0}-h_{k}(\mathbf{x}_{0})\|+T\frac{8mL_{h}\eta M}{|\mathcal{B}_{c}|\gamma_{2}^{1/2}}+T\frac{4\gamma_{2}^{1/2}\sigma_{h}}{|\mathcal{B}_{2,k}|^{1/2}}\Big{]},$
where we use the fact that $\sum_{t=0}^{T-1}(1-\mu)^{t}\leq\frac{1}{\mu}$ for all $\mu\in[0,1]$. Lower bounding the left-hand-side by $\min_{\mathbf{x}}\Phi(\mathbf{x})$ and dividing both sides by $T$, we obtain
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla\Phi_{1/\bar{\rho}}(\mathbf{{x}}_{t})\|^{2}]$
$\leq\frac{2}{\eta T}[\Phi_{1/\bar{\rho}}(\mathbf{x}_{0})-\min_{\mathbf{x}}\Phi(\mathbf{x})]+\eta\bar{\rho}M^{2}+\frac{\mathcal{C}}{T}\frac{m\beta\bar{\rho}}{\gamma_{2}|\mathcal{B}_{c}|}+\mathcal{C}\Big{(}\frac{m\bar{\rho}\beta\eta M}{|\mathcal{B}_{c}|\gamma_{2}^{1/2}}+\frac{\beta\bar{\rho}\gamma_{2}^{1/2}}{|\mathcal{B}_{2,k}|^{1/2}}\Big{)},$
where $\mathcal{C}=\max\{\frac{8}{m}\sum_{k=1}^{m}\|u_{2,k}^{0}-h_{k}(\mathbf{x}_{0})\|,32L_{h},16\sigma_{h}\}$. As we set $M^{2}\geq 2\sigma_{f}^{2}+2L_{F}^{2}+2\beta^{2}L_{h}^{2}$ and $\frac{\rho_{0}}{2}+\beta\frac{\rho_{1}}{2}+\frac{C}{2}=\frac{\bar{\rho}}{2}$, with $\gamma_{2}=\mathcal{O}(\frac{|\mathcal{B}_{2,k}|\epsilon^{4}}{\beta^{4}})$, $\eta=\mathcal{O}(\frac{|\mathcal{B}_{c}||\mathcal{B}_{2,k}|^{1/2}\epsilon^{4}}{\beta^{5}m})$ and $T=\mathcal{O}(\frac{\beta^{6}m}{|\mathcal{B}_{c}||\mathcal{B}_{2,k}|^{1/2}\epsilon^{6}})$, we have
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla\Phi_{1/\bar{\rho}}(\mathbf{{x}}_{t})\|^{2}]\leq\mathcal{O}(\epsilon^{2}).$
By Jensen’s inequality, we can get $\mathbb{E}[\|\nabla\Phi_{1/\bar{\rho}}(\mathbf{{x}}_{\hat{t}})\|]\leq\mathcal{O}(\epsilon)$, where $\hat{t}$ is selected uniformly at random from $\{1,\ldots,T\}$.
Appendix D Proof of Theorem 5.5 and D.1
First, we present the Algorithm 2 in this Appendix.
Theorem D.1.
Suppose Assumption 3.1, 5.1 and 5.4 with condition(ii) hold. Let $\gamma=\mathcal{O}(\min\{\frac{|\mathcal{B}_{2,k}|}{\beta^{4}}\epsilon^{4},\frac{|\mathcal{B}_{1,i}|}{\beta}\epsilon^{2}\})$, $\eta=\mathcal{O}(\min\{\frac{|\mathcal{B}||\mathcal{B}_{1,i}|^{1/2}\epsilon^{2}}{n\beta^{2}},\frac{|\mathcal{B}_{c}||\mathcal{B}_{2,k}|^{1/2}\epsilon^{4}}{\beta^{5}m}\})$. After $T=\mathcal{O}(\max\{\frac{m\beta^{6}}{|\mathcal{B}_{2,k}|^{1/2}|\mathcal{B}_{c}|\epsilon^{6}},\frac{n\beta^{3}}{|\mathcal{B}||\mathcal{B}_{1,i}|^{1/2}\epsilon^{4}},\frac{n\beta^{2}}{|\mathcal{B}||\mathcal{B}_{1,i}|\epsilon^{4}}\})$ iterations, the Algorithm 2 satisfies $\mathbb{E}[\|\nabla\Phi_{\lambda}(\mathbf{x}_{\hat{t}})\|]\leq\epsilon$ from some $\lambda=O(1/(\rho_{0}+\rho_{1}\beta))$, where $\hat{t}$ is selected uniformly at random from $\{1,\ldots,T\}$
Proof.
The objective function is $\min_{\mathbf{x}}\Phi(\mathbf{x},\mathcal{D}):=\underbrace{\frac{1}{n}\sum_{i=1}^{n}f_{i}(g_{i}(\mathbf{x}))}_{F(\mathbf{x})}+\underbrace{\frac{\beta}{m}\sum_{k=1}^{m}[h_{k}(\mathbf{x})]_{+}}_{H(\mathbf{x})}$.
To this end, we first give the proof of Theorem 5.5.
For simplicity, denote $\bar{\mathbf{x}}_{t}:=\text{prox}_{\Phi/\bar{\rho}}(\mathbf{x}_{t})$. Consider change in the Moreau envelope
Algorithm 2 Algorithm for solving (3) under Setting II
1:  Initialization: choose $\mathbf{x}_{0},\beta,\gamma_{1},\gamma_{2}$ and $\eta$.
2:  for $t=0$ to $T-1$ do
3:     Sample $\mathcal{B}^{t}\in\{1,\ldots,n\}$, $\mathcal{B}^{t}_{c}\in\{1,\ldots,m\}$
4:     for  each $i\in\mathcal{B}^{t}$  do
5:        Sample a mini-batch $\mathcal{B}_{1,i}^{t}$
6:        Update $u_{1,i}^{t}$ by Eqn.(15)
7:     end for
8:     Let $u_{1,i}^{t+1}=u_{1,i}^{t},i\notin\mathcal{B}^{t}$
9:     Compute $G_{1}^{t}$ in Eqn (16)
10:     for  each $k\in\mathcal{B}_{c}^{t}$ do
11:        Sample a mini-batch $\mathcal{B}_{2,k}^{t}$
12:        Update the estimator $u_{2,k}^{t+1}$ by Eqn (12)
13:     end for
14:     Let $u_{2,k}^{t+1}=u_{2,k}^{t},k\notin\mathcal{B}_{c}^{t}$
15:     Compute $G_{2}^{t}$ as in Eqn (13)
16:     Update $\mathbf{x}_{t+1}=\mathbf{x}_{t}-\eta(G_{1}^{t}+G_{2}^{t})$
17:  end for
$\mathbb{E}_{t}[\Phi_{1/\bar{\rho}}(\mathbf{x}_{t+1})]$ $=\mathbb{E}_{t}[\min_{\tilde{\mathbf{x}}}\Phi(\tilde{\mathbf{x}})+\frac{\bar{\rho}}{2}\|\tilde{\mathbf{x}}-\mathbf{x}_{t+1}\|^{2}]$
$\leq\mathbb{E}_{t}[\Phi(\bar{\mathbf{x}}_{t})+\frac{\bar{\rho}}{2}\|\bar{\mathbf{x}}_{t}-\mathbf{x}_{t+1}\|^{2}]$
$=\mathbb{E}_{t}[\Phi(\bar{\mathbf{x}}_{t})+\frac{\bar{\rho}}{2}\|\bar{\mathbf{x}}_{t}-(\mathbf{x}_{t}-\eta(G_{1}^{t}+G_{2}^{t}))\|^{2}]$
$\leq\Phi(\bar{\mathbf{x}}_{t})+\frac{\bar{\rho}}{2}\|\bar{\mathbf{x}}_{t}-\mathbf{x}_{t}\|^{2}+\bar{\rho}\mathbb{E}_{t}[\eta\langle\bar{\mathbf{x}}_{t}-\mathbf{x}_{t},G_{1}^{t}+G_{2}^{t}\rangle]+\eta^{2}\frac{\bar{\rho}M^{2}}{2}$
$=\Phi_{1/\bar{\rho}}(\mathbf{x}_{t})+\bar{\rho}\eta\langle\bar{\mathbf{x}}_{t}-\mathbf{x}_{t},\mathbb{E}_{t}[G_{1}^{t}]\rangle+\bar{\rho}\eta\langle\bar{\mathbf{x}}_{t}-\mathbf{x}_{t},\mathbb{E}_{t}[G_{2}^{t}]\rangle+\eta^{2}\frac{\bar{\rho}M^{2}}{2},$ (31)
where the second inequality uses the bound of $\mathbb{E}_{t}[\|G_{1}^{t}+G_{2}^{t}\|^{2}]$, which follows from the Lipchitz continuity, denoted by $M^{2}\geq 2L_{f}^{2}L_{g}^{2}+2\beta^{2}L_{h}^{2}$.
$\mathbb{E}_{t}[G_{1}^{t}]\in\frac{1}{n}\sum_{i=1}^{n}\partial f_{i}(u_{1,i}^{t})^{\top}\partial{g_{i}}(\mathbf{x}_{t})^{\top},~{}~{}~{}~{}\mathbb{E}_{t}[G_{2}^{t}]\in\frac{\beta}{m}\sum_{k=1}^{m}\partial h_{k}(\mathbf{x}_{t})[u_{2,k}^{t}]^{\prime}_{+}.$
Next we give the estimation of $\langle\bar{\mathbf{x}}_{t}-\mathbf{x}_{t},\mathbb{E}_{t}[G_{1}^{t}]\rangle$ and $\langle\bar{\mathbf{x}}_{t}-\mathbf{x}_{t},\mathbb{E}_{t}[G_{2}^{t}]\rangle$. For given $i\in\{1,\ldots,n\}$, we have
$f_{i}({g_{i}}(\bar{\mathbf{x}}_{t}))-f_{i}(u_{1,i}^{t})\geq\partial f_{i}(u_{1,i}^{t})^{\top}({g_{i}}(\bar{\mathbf{x}}_{t})-u_{1,i}^{t})-\frac{\rho_{f}}{2}\|{g_{i}}(\bar{\mathbf{x}}_{t})-u_{1,i}^{t}\|^{2}$
$\geq\partial f_{i}(u_{1,i}^{t})^{\top}({g_{i}}(\bar{\mathbf{x}}_{t})-u_{1,i}^{t})-\rho_{f}\|{g_{i}}(\bar{\mathbf{x}}_{t})-{g_{i}}(\mathbf{x}_{t})\|^{2}-\rho_{f}\|{g_{i}}(\mathbf{x}_{t})-u_{1,i}^{t}\|^{2}$
$\geq\partial f_{i}(u_{1,i}^{t})^{\top}\left[{g_{i}}(\mathbf{x}_{t})-u_{1,i}^{t}+\partial{g_{i}}(\mathbf{x}_{t})^{\top}(\bar{\mathbf{x}}_{t}-\mathbf{x}_{t})-\frac{\rho_{g}}{2}\|\bar{\mathbf{x}}_{t}-\mathbf{x}_{t}\|^{2}\right]-\rho_{f}L_{g}^{2}\|\bar{\mathbf{x}}_{t}-\mathbf{x}_{t}\|^{2}-\rho_{f}\|{g_{i}}(\mathbf{x}_{t})-u_{1,i}^{t}\|^{2}$
$\geq\partial f_{i}(u_{1,i}^{t})^{\top}({g_{i}}(\mathbf{x}_{t})-u_{1,i}^{t})+\partial f_{i}(u_{1,i}^{t})^{\top}\partial{g_{i}}(\mathbf{x}_{t})^{\top}(\bar{\mathbf{x}}_{t}-\mathbf{x}_{t})-(\frac{\rho_{g}L_{f}}{2}+\rho_{f}L_{g}^{2})\|\bar{\mathbf{x}}_{t}-\mathbf{x}_{t}\|^{2}-\rho_{f}\|{g_{i}}(\mathbf{x}_{t})-u_{1,i}^{t}\|^{2},$
where the first inequality holds by $\rho_{f}$-weakly convex of $f_{i}$, the second inequality holds by the monotonically non-decreasing of $f_{i}$ and the third inequality holds by $\rho_{g}$-weakly convexity of $g_{i}$, the last inequality holds by and $0\leq\partial f_{i}(u_{i}^{t})\leq L_{f}$. Then it follows
$\frac{1}{n}\sum_{i=1}^{n}\partial f_{i}(u_{1,i}^{t})^{\top}\partial{g_{i}}(\mathbf{x}_{t})^{\top}(\bar{\mathbf{x}}_{t}-\mathbf{x}_{t})$ $\leq\frac{1}{n}\sum_{i=1}^{n}\Big{[}f_{i}({g_{i}}(\bar{\mathbf{x}}_{t}))-f_{i}(u_{1,i}^{t})-\partial f_{i}(u_{1,i}^{t})^{\top}({g_{i}}(\mathbf{x}_{t})-u_{1,i}^{t})$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+(\frac{\rho_{g}L_{f}}{2}+\rho_{f}L_{g}^{2})\|\bar{\mathbf{x}}_{t}-\mathbf{x}_{t}\|^{2}+\rho_{f}\|{g_{i}}(\mathbf{x}_{t})-u_{1,i}^{t}\|^{2}\Big{]}.$ (32)
Next we bound $\langle\bar{\mathbf{x}}_{t}-\mathbf{x}_{t},\mathbb{E}_{t}[G_{2}^{t}]\rangle$. For given $k\in\{1,\ldots,m\}$, from (27), we have
$\beta\frac{1}{m}\sum_{k=1}^{m}[u_{2,k}^{t}]^{\prime}_{+}\partial{h_{k}}(\mathbf{x}_{t})^{\top}(\bar{\mathbf{x}}_{t}-\mathbf{x}_{t})$ $\leq\frac{\beta}{m}\sum_{k=1}^{m}\Big{[}[{h_{k}}(\bar{\mathbf{x}}_{t})]_{+}-[u_{2,k}^{t}]_{+}-[u_{2,k}^{t}]^{\prime}_{+}({h_{k}}(\mathbf{x}_{t})-u_{2,k}^{t})+\frac{\rho_{1}}{2}\|\bar{\mathbf{x}}_{t}-\mathbf{x}_{t}\|^{2}\Big{]}.$ (33)
Combining inequality (31), (32) and (33) yields
$\mathbb{E}_{t}[\Phi_{1/\bar{\rho}}(\mathbf{x}_{t+1})]\leq\Phi_{1/\bar{\rho}}({\mathbf{x}_{t}})+\bar{\rho}\eta\langle\bar{\mathbf{x}}_{t}-\mathbf{x}_{t},\mathbb{E}_{t}[G_{1}^{t}]\rangle+\bar{\rho}\eta\langle\bar{\mathbf{x}}_{t}-\mathbf{x}_{t},\mathbb{E}_{t}[G_{2}^{t}]\rangle+\eta^{2}\frac{\bar{\rho}M^{2}}{2}$
$\leq\Phi_{1/\bar{\rho}}({\mathbf{x}_{t}})+\eta^{2}\frac{\bar{\rho}M^{2}}{2}$
$+\bar{\rho}\eta\frac{1}{n}\sum_{i=1}^{n}\Big{[}f_{i}({g_{i}}(\bar{\mathbf{x}}_{t}))-f_{i}(u_{1,i}^{t})-\partial f_{i}(u_{1,i}^{t})^{\top}({g_{i}}(\mathbf{x}_{t})-u_{1,i}^{t})+(\frac{\rho_{g}L_{f}}{2}+\rho_{f}L_{g}^{2})\|\bar{\mathbf{x}}_{t}-\mathbf{x}_{t}\|^{2}+\rho_{f}\|{g_{i}}(\mathbf{x}_{t})-u_{1,i}^{t}\|^{2}\Big{]}$
$+\bar{\rho}\eta\frac{\beta}{m}\sum_{k=1}^{m}\Big{[}[{h_{k}}(\bar{\mathbf{x}}_{t})]_{+}-[u_{2,k}^{t}]_{+}-[u_{2,k}^{t}]^{\prime}_{+}({h_{k}}(\mathbf{x}_{t})-u_{2,k}^{t})+\frac{\rho_{1}}{2}\|\bar{\mathbf{x}}_{t}-\mathbf{x}_{t}\|^{2}\Big{]}$
$=\Phi_{1/\bar{\rho}}(\mathbf{x}_{t})+\eta^{2}\frac{\bar{\rho}M^{2}}{2}+\bar{\rho}\eta\frac{1}{n}\sum_{i=1}^{n}(\frac{\rho_{g}L_{f}}{2}+\rho_{f}L_{g}^{2})\|\bar{\mathbf{x}}_{t}-\mathbf{x}_{t}\|^{2}+\bar{\rho}\eta\frac{\beta}{m}\sum_{k=1}^{m}\frac{\rho_{1}}{2}\|\bar{\mathbf{x}}_{t}-\mathbf{x}_{t}\|^{2}$
$+\bar{\rho}\eta\frac{1}{n}\sum_{i=1}^{n}\Big{[}f_{i}({g_{i}}(\bar{\mathbf{x}}_{t}))-f_{i}(g_{i}(\mathbf{x}_{t}))+f_{i}(g_{i}(\mathbf{x}_{t}))-f_{i}(u_{1,i}^{t})-\partial f_{i}(u_{1,i}^{t})^{\top}({g_{i}}(\mathbf{x}_{t})-u_{i}^{t})+\rho_{f}\|{g_{i}}(\mathbf{x}_{t})-u_{1,i}^{t}\|^{2}\Big{]}$
$+\bar{\rho}\eta\frac{\beta}{m}\sum_{k=1}^{m}\Big{[}[{h_{k}}(\bar{\mathbf{x}}_{t})]_{+}-[{h_{k}}(\mathbf{x}_{t})]_{+}+[{h_{k}}(\mathbf{x}_{t})]_{+}-[u_{2,k}^{t}]_{+}-[u_{2,k}^{t}]^{\prime}_{+}({h_{k}}(\mathbf{x}_{t})-u_{2,k}^{t})\Big{]}.$ (34)
From Lemma 4.1, the function $\Phi$ is $C$-weakly convex. We have $(\bar{\rho}-C)-$strong convexity of $\mathbf{x}\mapsto\Phi(\mathbf{x})+\frac{\bar{\rho}}{2}\|\mathbf{x}_{t}-\mathbf{x}\|^{2}$
$\Phi(\mathbf{\bar{x}}_{t})-\Phi(\mathbf{x}_{t})\leq(\frac{C}{2}-\bar{\rho})\|\mathbf{\bar{x}}_{t}-\mathbf{x}_{t}\|^{2}.$
We get
$\frac{1}{n}\sum_{i=1}^{n}f_{i}({g_{i}}(\bar{\mathbf{x}}_{t}))-f_{i}(g_{i}(\mathbf{x}_{t}))+\beta\frac{1}{m}\sum_{k=1}^{m}([{h_{k}}(\bar{\mathbf{x}}_{t})]_{+}-[{h_{k}}(\mathbf{x}_{t})]_{+})=\Phi(\mathbf{\bar{x}}_{t})-\Phi(\mathbf{x}_{t})\leq(\frac{C}{2}-\bar{\rho})\|\mathbf{\bar{x}}_{t}-\mathbf{x}_{t}\|^{2}.$ (35)
Plugging inequalities (35) to (34), we get
$\mathbb{E}_{t}[\Phi_{1/\bar{\rho}}(\mathbf{x}_{t+1})]$
$\leq\Phi_{1/\bar{\rho}}(\mathbf{x}_{t})+\eta^{2}\frac{\bar{\rho}M^{2}}{2}+\bar{\rho}\eta\Big{(}\frac{\rho_{g}L_{f}}{2}+\rho_{f}L_{g}^{2}+\beta\frac{\rho_{1}}{2}+\frac{C}{2}-\bar{\rho}\Big{)}\|\bar{\mathbf{x}}_{t}-\mathbf{x}_{t}\|^{2}$
$+\bar{\rho}\eta\frac{1}{n}\sum_{i=1}^{n}\Big{[}f_{i}(g_{i}(\mathbf{x}_{t}))-f_{i}(u_{1,i}^{t}))-\partial f_{i}(u_{1,i}^{t})^{\top}({g_{i}}(\mathbf{x}_{t})-u_{1,i}^{t})+\rho_{f}\|{g_{i}}(\mathbf{x}_{t})-u_{1,i}^{t}\|^{2}\Big{]}$
$+\bar{\rho}\eta\frac{\beta}{m}\sum_{k=1}^{m}\Big{[}[{h_{k}}(\mathbf{x}_{t})]_{+}-[u_{2,k}^{t}]_{+}-[u_{2,k}^{t}]^{\prime}_{+}({h_{k}}(\mathbf{x}_{t})-u_{2,k}^{t})\Big{]}.$
Set $\frac{\rho_{g}L_{f}}{2}+\rho_{f}L_{g}^{2}+\beta\frac{\rho_{1}}{2}+\frac{C}{2}=\frac{\bar{\rho}}{2}$, we have
$\mathbb{E}_{t}[\Phi_{1/\bar{\rho}}(\mathbf{x}_{t+1})]$
$\leq\Phi_{1/\bar{\rho}}(\mathbf{x}_{t})+\eta^{2}\frac{\bar{\rho}M^{2}}{2}-\eta\frac{\bar{\rho}^{2}}{2}\|\bar{\mathbf{x}}_{t}-\mathbf{x}_{t}\|^{2}$
$+\bar{\rho}\eta\frac{1}{n}\sum_{i=1}^{n}\Big{[}f_{i}(g_{i}(\mathbf{x}_{t}))-f_{i}(u_{1,i}^{t}))-\partial f_{i}(u_{1,i}^{t})^{\top}({g_{i}}(\mathbf{x}_{t})-u_{1,i}^{t})+\rho_{f}\|{g_{i}}(\mathbf{x}_{t})-u_{1,i}^{t}\|^{2}\Big{]}$
$+\bar{\rho}\eta\frac{\beta}{m}\sum_{k=1}^{m}\Big{[}[{h_{k}}(\mathbf{x}_{t})]_{+}-[u_{2,k}^{t}]_{+}-[u_{2,k}^{t}]^{\prime}_{+}({h_{k}}(\mathbf{x}_{t})-u_{2,k}^{t})\Big{]}$
$\leq\Phi_{1/\bar{\rho}}(\mathbf{x}_{t})+\eta^{2}\frac{\bar{\rho}M^{2}}{2}-\frac{\eta}{2}\|\nabla\Phi_{1/\bar{\rho}}(\mathbf{x}_{t})\|^{2}$
$+\bar{\rho}\eta\frac{1}{n}\sum_{i=1}^{n}\Big{[}f_{i}(g_{i}(\mathbf{x}_{t}))-f_{i}(u_{1,i}^{t}))-\partial f_{i}(u_{1,i}^{t})^{\top}({g_{i}}(\mathbf{x}_{t})-u_{1,i}^{t})+\rho_{f}\|{g_{i}}(\mathbf{x}_{t})-u_{1,i}^{t}\|^{2}\Big{]}$
$+\bar{\rho}\eta\frac{\beta}{m}\sum_{k=1}^{m}\Big{[}[{h_{k}}(\mathbf{x}_{t})]_{+}-[u_{2,k}^{t}]_{+}-[u_{2,k}^{t}]^{\prime}_{+}({h_{k}}(\mathbf{x}_{t})-u_{2,k}^{t})\Big{]}.$
Using the Lipschitz continuity of $f_{i}$ and the fact that $0\leq[u_{2,k}^{t}]^{\prime}_{+}\leq 1$, we get
$\mathbb{E}_{t}[\Phi_{1/\bar{\rho}}(\mathbf{x}_{t+1})]$ $\leq\Phi_{1/\bar{\rho}}(\mathbf{x}_{t})+\eta^{2}\frac{\bar{\rho}M^{2}}{2}-\frac{\eta}{2}\|\nabla\Phi_{1/\bar{\rho}}(\mathbf{x}_{t})\|^{2}+2\bar{\rho}L_{f}\eta\frac{1}{n}\sum_{i=1}^{n}\|g_{i}(\mathbf{x}_{t})-u_{1,i}^{t}\|$
$+\bar{\rho}\rho_{f}\eta\frac{1}{n}\sum_{i=1}^{n}\|{g_{i}}(\mathbf{x}_{t})-u_{1,i}^{t}\|^{2}+2\bar{\rho}\eta\frac{\beta}{m}\sum_{k=1}^{m}\|{h_{k}}(\mathbf{x}_{t})-u_{2,k}^{t}\|.$ (36)
Taking the full expectation on both sides of (36) and adding the result of Lemma B.1, (21) and (20), we have
$\mathbb{E}[\Phi_{1/\bar{\rho}}(\mathbf{x}_{t+1})]$
$\leq\mathbb{E}[\Phi_{1/\bar{\rho}}(\mathbf{x}_{t})]+\eta^{2}\frac{\bar{\rho}M^{2}}{2}-\frac{\eta}{2}\mathbb{E}[\|\nabla\Phi_{1/\bar{\rho}}(\mathbf{x}_{t})\|^{2}]$
$+2\bar{\rho}\eta\Big{[}L_{f}(1-\frac{|\mathcal{B}|\gamma_{1}}{2n})^{t}\frac{1}{n}\sum_{i=1}^{n}\|u_{1,i}^{0}-g_{i}(\mathbf{x}_{0})\|+\frac{4nL_{g}L_{f}M\eta}{|\mathcal{B}|\gamma_{1}^{1/2}}+\frac{2L_{f}\gamma_{1}^{1/2}\sigma_{g}}{|\mathcal{B}_{1,i}|^{1/2}}\Big{]}$
$+\bar{\rho}\eta\Big{[}\rho_{f}(1-\frac{|\mathcal{B}|\gamma_{1}}{n})^{t}\frac{1}{n}\sum_{i=1}^{n}\|u_{1,i}^{0}-g_{i}(\mathbf{x}_{0})\|^{2}+\frac{8n^{2}\eta^{2}L_{g}^{2}\rho_{f}M^{2}}{|\mathcal{B}|^{2}\gamma_{1}}+\frac{2\gamma_{1}\sigma_{g}^{2}\rho_{f}}{|\mathcal{B}_{1,i}|}\Big{]}$
$+2\bar{\rho}\eta\beta\Big{[}(1-\frac{|\mathcal{B}_{c}|\gamma_{2}}{2m})^{t}\frac{1}{m}\sum_{k=1}^{m}\|u_{2,k}^{0}-h_{k}(\mathbf{x}_{0})\|+\frac{4mL_{h}\eta M}{|\mathcal{B}_{c}|\gamma_{2}^{1/2}}+\frac{2\gamma_{2}^{1/2}\sigma_{h}}{|\mathcal{B}_{2,k}|^{1/2}}\Big{]}.$
Taking summation from $t=0$ to $T-1$ yields
$\mathbb{E}[\Phi_{1/\bar{\rho}}(\mathbf{x}_{T})]$
$\leq\Phi_{1/\bar{\rho}}(\mathbf{x}_{0})+\eta^{2}T\frac{\bar{\rho}M^{2}}{2}-\frac{\eta}{2}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla\Phi_{1/\bar{\rho}}(\mathbf{x}_{t})\|^{2}]$
$+2\bar{\rho}\eta\Big{[}L_{f}\frac{2n}{|\mathcal{B}|\gamma_{1}}\frac{1}{n}\sum_{i=1}^{n}\|u_{1,i}^{0}-g_{i}(\mathbf{x}_{0})\|+T\frac{8nL_{g}L_{f}\eta M}{|\mathcal{B}|\gamma_{1}^{1/2}}+T\frac{4L_{f}\gamma_{1}^{1/2}\sigma_{g}}{|\mathcal{B}_{1,i}|^{1/2}}\Big{]}$
$+\bar{\rho}\eta\Big{[}\rho_{f}\frac{n}{|\mathcal{B}|\gamma_{1}}\frac{1}{n}\sum_{i=1}^{n}\|u_{1,i}^{0}-g_{i}(\mathbf{x}_{0})\|^{2}+T\frac{8n^{2}\eta^{2}L_{g}^{2}\rho_{f}M^{2}}{|\mathcal{B}|^{2}\gamma_{1}}+T\frac{2\gamma_{1}\sigma_{g}^{2}\rho_{f}}{|\mathcal{B}_{1,i}|}\Big{]}$
$+2\bar{\rho}\eta\beta\Big{[}\frac{2m}{|\mathcal{B}_{c}|\gamma_{2}}\frac{1}{m}\sum_{k=1}^{m}\|u_{2,k}^{0}-h_{k}(\mathbf{x}_{0})\|+T\frac{8mL_{h}\eta M}{|\mathcal{B}_{c}|\gamma_{2}^{1/2}}+T\frac{4\gamma_{2}^{1/2}\sigma_{h}}{|\mathcal{B}_{2,k}|^{1/2}}\Big{]},$
where we use the fact that $\sum_{t=0}^{T-1}(1-\mu)\leq\frac{1}{\mu}$. Lower bounding the left-hand-side by $\min_{\mathbf{x}}\Phi(\mathbf{x})$ and dividing both sides by $T$, we obtain
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla\Phi_{1/\bar{\rho}}(\mathbf{{w}}_{t})\|^{2}]$
$\leq\frac{2}{\eta T}[\Phi_{1/\bar{\rho}}(\mathbf{x}_{0})-\min_{\mathbf{x}}\Phi(\mathbf{x})]+\eta\bar{\rho}M^{2}+\frac{\mathcal{C}\bar{\rho}}{T\gamma}\max\{\frac{n}{|\mathcal{B}|},\frac{\beta m}{|\mathcal{B}_{c}|}\}$
$+\mathcal{C}\Big{(}\bar{\rho}\frac{\eta M}{\gamma^{1/2}}\max\{\frac{n}{|\mathcal{B}|},\frac{\beta m}{|\mathcal{B}_{c}|}\}+\frac{n^{2}\eta^{2}M^{2}\bar{\rho}}{|\mathcal{B}|^{2}\gamma}+\gamma^{1/2}\bar{\rho}\max\{\frac{1}{|\mathcal{B}_{1,i}|^{1/2}},\frac{\beta}{|\mathcal{B}_{2,k}|^{1/2}}\}+\bar{\rho}\frac{\gamma}{|\mathcal{B}_{1,i}|}\Big{)}.$
where we set $\gamma_{1}=\gamma_{2}=\gamma$ and
[…]
$32L_{g}L_{f},16L_{g}^{2}\rho_{f},16L_{f}\sigma_{g},4\sigma_{g}^{2}\rho_{f},16L_{h},16\sigma_{h}\}.$
As we set $M^{2}\geq 2L_{f}^{2}L_{g}^{2}+2\beta^{2}L_{h}^{2}$ and $\frac{\rho_{g}L_{f}}{2}+\rho_{f}L_{g}^{2}+\beta\frac{\rho_{1}}{2}+\frac{C}{2}=\frac{\bar{\rho}}{2}$, with $\gamma=\mathcal{O}(\min\{|\mathcal{B}_{1,i}|,\frac{|\mathcal{B}_{2,k}|}{\beta^{2}}\}\frac{\epsilon^{4}}{\beta^{2}})$, $\eta=\mathcal{O}(\min\{\frac{|\mathcal{B}|}{n},\frac{|\mathcal{B}_{c}|}{\beta m}\}\min\{|\mathcal{B}_{1,i}|^{1/2},\frac{|\mathcal{B}_{2,k}|^{1/2}}{\beta}\}\frac{\epsilon^{4}}{\beta^{3}})$ and $T=\mathcal{O}(\max\{\frac{\beta}{|\mathcal{B}_{1,i}|^{1/2}},\frac{\beta^{2}}{|\mathcal{B}_{2,k}|^{1/2}},\frac{1}{|\mathcal{B}_{1,i}|}\}\max\{\frac{n}{|\mathcal{B}|},\frac{\beta m}{|\mathcal{B}_{c}|}\}\frac{\beta^{3}}{\epsilon^{6}})$, we have
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla\Phi_{1/\bar{\rho}}(\mathbf{{x}}_{t})\|^{2}]\leq\mathcal{O}(\epsilon^{2}).$
By Jensen’s inequality, we can get $\mathbb{E}[\|\nabla\Phi_{1/\bar{\rho}}(\mathbf{{x}}_{\hat{t}})\|]\leq\mathcal{O}(\epsilon)$, where $\hat{t}$ is selected uniformly at random from $\{1,\ldots,T\}$.
Now, we give the proof of Theorem D.1.
From (31), also denote $\bar{\mathbf{x}}_{t}:=\text{prox}_{\Phi/\bar{\rho}}(\mathbf{x}_{t})$ and consider change in the Moreau envelope, we have
$\mathbb{E}_{t}[\Phi_{1/\bar{\rho}}(\mathbf{x}_{t+1})]\leq\Phi_{1/\bar{\rho}}(\mathbf{x}_{t})+\bar{\rho}\eta\langle\bar{\mathbf{x}}_{t}-\mathbf{x}_{t},\mathbb{E}_{t}[G_{1}^{t}]\rangle+\bar{\rho}\eta\langle\bar{\mathbf{x}}_{t}-\mathbf{x}_{t},\mathbb{E}_{t}[G_{2}^{t}]\rangle+\eta^{2}\frac{\bar{\rho}M^{2}}{2},$ (37)
where the gradient can be bounded by $M$, where $M^{2}\geq 2L_{f}^{2}L_{g}^{2}+2\beta^{2}L_{h}^{2}$ and
$\mathbb{E}_{t}[G_{1}^{t}]\in\frac{1}{n}\sum_{i=1}^{n}\nabla f_{i}(u_{1,i}^{t})^{\top}\nabla{g_{i}}(\mathbf{x}_{t})^{\top},~{}~{}~{}~{}\mathbb{E}_{t}[G_{2}^{t}]\in\frac{\beta}{m}\sum_{k=1}^{m}\partial h_{k}(\mathbf{x}_{t})[u_{2,k}^{t}]^{\prime}_{+}.$
Next we give the bound of $\langle\bar{\mathbf{x}}_{t}-\mathbf{x}_{t},\mathbb{E}_{t}[G_{2}^{t}]\rangle$ and $\langle\bar{\mathbf{x}}_{t}-\mathbf{x}_{t},\mathbb{E}_{t}[G_{2}^{t}]\rangle$. It’s easy to verify that $F$ is $L_{\nabla F}$-smooth, where $L_{\nabla F}=L_{f}L_{\nabla g}+L_{g}^{2}L_{\nabla f}$. Then we have
$F(\mathbf{x}_{t})\leq F(\bar{\mathbf{x}}_{t})+\langle\nabla F(\bar{\mathbf{x}}_{t}),\mathbf{x}^{t}-\bar{\mathbf{x}}_{t}\rangle+\frac{L_{\nabla F}}{2}\|\bar{\mathbf{x}}_{t}-\mathbf{x}_{t}\|^{2}$
$=F(\bar{\mathbf{x}}_{t})+\langle\nabla F(\bar{\mathbf{x}}_{t})-\nabla F(\mathbf{x}_{t}),\mathbf{x}_{t}-\bar{\mathbf{x}}_{t}\rangle+\frac{L_{\nabla F}}{2}\|\bar{\mathbf{x}}_{t}-\mathbf{x}_{t}\|^{2}$
$+\langle\frac{1}{n}\sum_{i=1}^{n}\nabla f_{i}({g_{i}}(\mathbf{x}_{t}))^{\top}\nabla{g_{i}}(\mathbf{x}_{t})^{\top}-\frac{1}{n}\sum_{i=1}^{n}\nabla f_{i}(u_{1,i}^{t})^{\top}\nabla{g_{i}}(\mathbf{x}_{t})^{\top},\mathbf{x}_{t}-\bar{\mathbf{x}}_{t}\rangle$
$+\langle\frac{1}{n}\sum_{i=1}^{n}\nabla f_{i}(u_{1,i}^{t})^{\top}\nabla{g_{i}}(\mathbf{x}_{t})^{\top},\mathbf{x}_{t}-\bar{\mathbf{x}}_{t}\rangle.$
Then, we have
$\mathbb{E}_{t}[\langle G_{1}^{t},\bar{\mathbf{x}}_{t}-\mathbf{x}_{t}\rangle]=\langle\frac{1}{n}\sum_{i=1}^{n}\nabla f_{i}(u_{1,i}^{t})^{\top}\nabla{g_{i}}(\mathbf{x}_{t})^{\top},\bar{\mathbf{x}}_{t}-\mathbf{x}_{t}\rangle$
$\leq F(\bar{\mathbf{x}}_{t})-F(\mathbf{x}_{t})+\frac{L_{\nabla F}}{2}\|\bar{\mathbf{x}}_{t}-\mathbf{x}_{t}\|^{2}+L_{\nabla F}\|\bar{\mathbf{x}}_{t}-\mathbf{x}_{t}\|^{2}+L_{\nabla f}L_{g}\frac{1}{n}\sum_{i=1}^{n}\|g_{i}(\mathbf{x}_{t})-u_{1,i}^{t}\|\|\bar{\mathbf{x}}_{t}-\mathbf{x}_{t}\|$
$\leq F(\bar{\mathbf{x}}_{t})-F(\mathbf{x}_{t})+(2L_{\nabla F}+2L_{\nabla f}L_{g})\|\bar{\mathbf{x}}_{t}-\mathbf{x}_{t}\|^{2}+2L_{\nabla f}L_{g}\frac{1}{n}\sum_{i=1}^{n}\|g_{i}(\mathbf{x}_{t})-u_{1,i}^{t}\|^{2}.$
From the (27), we have
$\mathbb{E}_{t}[\langle G_{2}^{t},\bar{\mathbf{x}}_{t}-\mathbf{x}_{t}\rangle]$ $=\beta\frac{1}{m}\sum_{k=1}^{m}[u_{2,k}^{t}]^{\prime}_{+}\partial{h_{k}}(\mathbf{x}_{t})^{\top}(\bar{\mathbf{x}}_{t}-\mathbf{x}_{t})$
$\leq\frac{\beta}{m}\sum_{k=1}^{m}\Big{[}[{h_{k}}(\bar{\mathbf{x}}_{t})]_{+}-[u_{2,k}^{t}]_{+}-[u_{2,k}^{t}]^{\prime}_{+}({h_{k}}(\mathbf{x}_{t})-u_{2,k}^{t})+\frac{\rho_{1}}{2}\|\bar{\mathbf{x}}_{t}-\mathbf{x}_{t}\|^{2}\Big{]}.$
Adding above two estimation to (37), we have
$\mathbb{E}_{t}[\Phi_{1/\bar{\rho}}(\mathbf{x}_{t+1})]\leq\Phi_{1/\bar{\rho}}(\mathbf{x}_{t})+\bar{\rho}\eta\langle\bar{\mathbf{x}}_{t}-\mathbf{x}_{t},\mathbb{E}_{t}[G_{1}^{t}]\rangle+\bar{\rho}\eta\langle\bar{\mathbf{x}}_{t}-\mathbf{x}_{t},\mathbb{E}_{t}[G_{2}^{t}]\rangle+\eta^{2}\frac{\bar{\rho}M^{2}}{2}$
$\leq\Phi_{1/\bar{\rho}}(\mathbf{x}_{t})+\eta^{2}\frac{\bar{\rho}M^{2}}{2}$
$+\bar{\rho}\eta\Big{[}F(\bar{\mathbf{x}}_{t})-F(\mathbf{x}_{t})+(2L_{\nabla F}+2L_{\nabla f}L_{g})\|\bar{\mathbf{x}}_{t}-\mathbf{x}_{t}\|^{2}+2L_{\nabla f}L_{g}\frac{1}{n}\sum_{i=1}^{n}\|g_{i}(\mathbf{x}_{t})-u_{1,i}^{t}\|^{2}\Big{]}$
$+\bar{\rho}\eta\frac{\beta}{m}\sum_{k=1}^{m}\Big{[}[{h_{k}}(\bar{\mathbf{x}}_{t})]_{+}-[{h_{k}}(\mathbf{x}_{t})]_{+}+[{h_{k}}(\mathbf{x}_{t})]_{+}-[u_{2,k}^{t}]_{+}-[u_{2,k}^{t}]^{\prime}_{+}({h_{k}}(\mathbf{x}_{t})-u_{2,k}^{t})+\frac{\rho_{1}}{2}\|\bar{\mathbf{x}}_{t}-\mathbf{x}_{t}\|^{2}\Big{]}.$ (38)
Since smoothness of $F$ implies the weakly convexity of $F$, we still have $(\bar{\rho}-C)-$strong convexity of $\mathbf{x}\mapsto\Phi(\mathbf{x})+\frac{\bar{\rho}}{2}\|\mathbf{x}_{t}-\mathbf{x}\|^{2}$
$\Phi(\mathbf{\bar{x}}_{t})-\Phi(\mathbf{x}_{t})\leq(\frac{C}{2}-\bar{\rho})\|\mathbf{\bar{x}}_{t}-\mathbf{x}_{t}\|^{2}.$
Adding above inequality to (38), we have
$\mathbb{E}_{t}[\Phi_{1/\bar{\rho}}(\mathbf{x}_{t+1})]$ $\leq\Phi_{1/\bar{\rho}}(\mathbf{x}_{t})+\eta^{2}\frac{\bar{\rho}M^{2}}{2}+\bar{\rho}\eta(2L_{\nabla F}+2L_{\nabla f}L_{g}+\beta\frac{\rho_{1}}{2}+\frac{C}{2}-\bar{\rho})\|\bar{\mathbf{x}}_{t}-\mathbf{x}_{t}\|^{2}$
$+2L_{\nabla f}L_{g}\bar{\rho}\eta\frac{1}{n}\sum_{i=1}^{n}\|g_{i}(\mathbf{x}_{t})-u_{1i}^{t}\|^{2}+\bar{\rho}\eta\frac{\beta}{m}\sum_{k=1}^{m}\Big{[}[{h_{k}}(\mathbf{x}_{t})]_{+}-[u_{2,k}^{t}]_{+}-[u_{2,k}^{t}]^{\prime}_{+}({h_{k}}(\mathbf{x}_{t})-u_{2,k}^{t})\Big{]}.$
If we set $2L_{\nabla F}+2L_{\nabla f}L_{f}+\beta\frac{\rho_{1}}{2}+\frac{C}{2}=\frac{\bar{\rho}}{2}$, we have
$\mathbb{E}_{t}[\Phi_{1/\bar{\rho}}(\mathbf{x}_{t+1})]$ $\leq\Phi_{1/\bar{\rho}}(\mathbf{x}_{t})+\eta^{2}\frac{\bar{\rho}M^{2}}{2}-\bar{\rho}\eta\frac{\bar{\rho}}{2}\|\bar{\mathbf{x}}_{t}-\mathbf{x}_{t}\|^{2}$
$+2L_{\nabla f}L_{g}\bar{\rho}\eta\frac{1}{n}\sum_{i=1}^{n}\|g_{i}(\mathbf{x}_{t})-u_{1i}^{t}\|^{2}+\bar{\rho}\eta\frac{\beta}{m}\sum_{k=1}^{m}\Big{[}[{h_{k}}(\mathbf{x}_{t})]_{+}-[u_{2,k}^{t}]_{+}-[u_{2,k}^{t}]^{\prime}_{+}({h_{k}}(\mathbf{x}_{t})-u_{2,k}^{t})\Big{]}$
$\leq\Phi_{1/\bar{\rho}}(\mathbf{x}_{t})+\eta^{2}\frac{\bar{\rho}M^{2}}{2}-\frac{\eta}{2}\|\nabla\Phi_{1/\bar{\rho}}(\mathbf{x}_{t})\|^{2}+2L_{\nabla f}\bar{\rho}\eta\frac{1}{n}\sum_{i=1}^{n}\|g_{i}(\mathbf{x}_{t})-u_{1i}^{t}\|^{2}+2\bar{\rho}\eta\frac{\beta}{m}\sum_{k=1}^{m}\|{h_{k}}(\mathbf{x}_{t})-u_{2,k}^{t}\|,$
where we use the fact $0\leq[u_{2,k}^{t}]^{\prime}_{+}\leq 1$.
Adding the result of Lemma B.1 and inequality (20), we get
$\mathbb{E}[\Phi_{1/\bar{\rho}}(\mathbf{x}_{t+1})]$ $\leq\mathbb{E}[\Phi_{1/\bar{\rho}}(\bar{\mathbf{x}}_{t})]+\eta^{2}\frac{\bar{\rho}M^{2}}{2}-\frac{\eta}{2}\mathbb{E}[\|\nabla\Phi_{1/\bar{\rho}}(\mathbf{x}_{t})\|^{2}]$
$+2\bar{\rho}\eta\Big{[}L_{\nabla f}L_{g}(1-\frac{|\mathcal{B}|\gamma_{1}}{n})^{t}\frac{1}{n}\sum_{i=1}^{n}\|u_{1,i}^{0}-g_{i}(\mathbf{x}_{0})\|^{2}+\frac{8n^{2}\eta^{2}L_{g}^{2}L_{\nabla f}L_{g}M^{2}}{|\mathcal{B}|^{2}\gamma_{1}}+\frac{2\gamma_{1}\sigma_{g}^{2}L_{\nabla f}L_{g}}{|\mathcal{B}_{1,i}|}\Big{]}$
$+2\bar{\rho}\eta\beta\Big{[}(1-\frac{|\mathcal{B}_{c}|\gamma_{2}}{2m})^{t}\frac{1}{m}\sum_{k=1}^{m}\|u_{2,k}^{0}-h_{k}(\mathbf{x}_{0})\|+\frac{4mL_{h}^{2}\eta M}{|\mathcal{B}_{c}|\gamma_{2}^{1/2}}+\frac{2\gamma_{2}^{1/2}\sigma_{h}}{|\mathcal{B}_{2,k}|^{1/2}}\Big{]}.$
Taking summation from $t=0$ to $T-1$ yields
$\mathbb{E}[\Phi_{1/\bar{\rho}}(\mathbf{x}_{T})]$
$\leq\mathbb{E}[\Phi_{1/\bar{\rho}}(\mathbf{x}_{0})]+\eta^{2}T\frac{\bar{\rho}M^{2}}{2}-\frac{\eta}{2}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla\Phi_{1/\bar{\rho}}(\mathbf{x}_{t})\|^{2}]$
$+2\bar{\rho}\eta\Big{[}L_{\nabla f}L_{g}\frac{n}{|\mathcal{B}|\gamma_{1}}\frac{1}{n}\sum_{i=1}^{n}\|u_{1,i}^{0}-g_{i}(\mathbf{x}_{0})\|^{2}+T\frac{8n^{2}\eta^{2}L_{g}^{3}L_{\nabla f}M^{2}}{|\mathcal{B}|^{2}\gamma_{1}}+T\frac{2\gamma_{1}\sigma_{g}^{2}L_{\nabla f}L_{g}}{|\mathcal{B}_{1,i}|}\Big{]}$
$+2\bar{\rho}\eta\beta\Big{[}\frac{2m}{|\mathcal{B}_{c}|\gamma_{2}}\frac{1}{m}\sum_{k=1}^{m}\|u_{2,k}^{0}-h_{k}(\mathbf{x}_{0})\|+T\frac{8mL_{h}\eta M}{|\mathcal{B}_{c}|\gamma_{2}^{1/2}}+T\frac{4\gamma_{2}^{1/2}\sigma_{h}}{|\mathcal{B}_{2,k}|^{1/2}}\Big{]},$
where we use the fact that $\sum_{t=0}^{T-1}(1-\mu)\leq\frac{1}{\mu}$.
Lower bounding the left-hand-side by $\min_{\mathbf{x}}\Phi(\mathbf{x})$ and dividing both sides by $T$, we obtain
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla\Phi_{1/\bar{\rho}}(\mathbf{{w}}_{t})\|^{2}]$
$\leq\frac{2}{\eta T}[\Phi_{1/\bar{\rho}}(\mathbf{x}_{0})-\min_{\mathbf{x}}\Phi(\mathbf{x})]+\eta\bar{\rho}M^{2}+\frac{\mathcal{C}\bar{\rho}}{T\gamma}\max\{\frac{n}{|\mathcal{B}|},\frac{\beta m}{|\mathcal{B}_{c}|}\}$
$+\mathcal{C}\Big{(}\frac{m\beta\eta\bar{\rho}M}{|\mathcal{B}_{c}|\gamma^{1/2}}+\frac{n^{2}\eta^{2}M^{2}\bar{\rho}}{|\mathcal{B}|^{2}\gamma}+\frac{\beta\bar{\rho}\gamma^{1/2}}{|\mathcal{B}_{2,k}|^{1/2}}+\frac{\gamma\bar{\rho}}{|\mathcal{B}_{1,i}|}\Big{)},$
where we use $\gamma_{1}=\gamma_{2}=\gamma$ and
[…]
$32L_{\nabla f}L_{g}^{3},32L_{h},8L_{\nabla f}L_{g}\sigma_{g}^{2},16\sigma_{h}\}.$
As we set $M^{2}\geq 2L_{f}^{2}L_{g}^{2}+2\beta^{2}L_{h}^{2}$ and $2L_{\nabla F}+2L_{\nabla f}L_{f}+\beta\frac{\rho_{1}}{2}+\frac{C}{2}=\frac{\bar{\rho}}{2}$, then with $\gamma=\mathcal{O}(\min\{\frac{|\mathcal{B}_{2,k}|}{\beta^{4}}\epsilon^{4},\frac{|\mathcal{B}_{1,i}|}{\beta}\epsilon^{2}\})$, $\eta=\mathcal{O}(\min\{\frac{|\mathcal{B}||\mathcal{B}_{1,i}|^{1/2}\epsilon^{2}}{n\beta^{2}},\frac{|\mathcal{B}_{c}||\mathcal{B}_{2,k}|^{1/2}\epsilon^{4}}{\beta^{5}m}\})$ and $T=\mathcal{O}(\max\{\frac{m\beta^{6}}{|\mathcal{B}_{2,k}|^{1/2}|\mathcal{B}_{c}|\epsilon^{6}},\frac{n\beta^{3}}{|\mathcal{B}||\mathcal{B}_{1,i}|^{1/2}\epsilon^{4}},\frac{n\beta^{2}}{|\mathcal{B}||\mathcal{B}_{1,i}|\epsilon^{4}}\})$, we have
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla\Phi_{1/\bar{\rho}}(\mathbf{{x}}_{t})\|^{2}]\leq\mathcal{O}(\epsilon^{2}).$
By Jensen’s inequality, we can get $\mathbb{E}[\|\nabla\Phi_{1/\bar{\rho}}(\mathbf{{x}}_{\hat{t}})\|]\leq\mathcal{O}(\epsilon)$, where $\hat{t}$ is selected uniformly at random from $\{1,\ldots,T\}$.